### v1

# Symplectic Lattices and GKP Codes - Simple Randomized Constructions from Cryptographic Lattices

Abstract
We construct good GKP (Gottesman-Kitaev-Preskill) codes (in the sense of Conrad, Eisert and Seifert proposed) from standard short integer solution (SIS) lattices as well as from ring SIS and module SIS lattices, R-SIS and M-SIS lattices, respectively. These lattice are crucial for lattice-based cryptography. Our construction yields GKP codes with distance $\sqrt{n/\pi e}$. This compares favorably with the NTRU-based construction by Conrad et al. that achieves distance $\Omega(\sqrt{n/q}),$ with $n\leq q^{2}/0.28$. Unlike their codes, our codes do not have secret keys that can be used to speed-up the decoding. However, we present a simple decoding algorithm that, for many parameter choices, experimentally yields decoding results similar to the ones for NTRU-based codes. Using the R-SIS and M-SIS construction, our simple decoding algorithm runs in nearly linear time. Following Conrad, Eisert and Seifert’s work, our construction of GKP codes follows directly from an explicit, randomized construction of symplectic lattices with (up to constants $\approx 1$) minimal distance $(1/\sigma_{2n})^{1/2n}\approx\sqrt{\frac{n}{\pi e}}$, where $\sigma_{2n}$ is the volume of the $2n$-dimensional unit ball. Before this result, Buser and Sarnak gave a non-constructive proof for the existence of such symplectic lattices.
Keywords: quantum error correction, good GKP codes, continuous-variable systems, symplectic lattices, minimal distances, short integer solutions
1 Introduction
Quantum error correction and fault tolerant quantum computing are essential ingredients for the realization of fully fledged quantum computing in the future. Which quantum error correction codes can be realized efficiently determines the kind of physical platforms most promising in building quantum computers; they influence the complete quantum computing stack, from the physical qubit level to the quantum software level. Bosonic quantum error-correcting codes encode discrete logical information in infinite-dimensional systems and promise hardware-efficient routes to fault tolerance on continuous-variable platforms, e. g. photonic-based computing platforms [1, 2, 3, 4]. A leading example are the Gottesman–Kitaev–Preskill (GKP) codes, stabilizer error correcting codes which place stabilizers on a regular grid in phase space so that small displacement errors can be detected and corrected. This idea originates with the seminal paper [5] by Gottesman, Kitaev and Preskill. Experimentally, grid encodings and error-correction primitives have been realized in superconducting cavities and trapped ions, demonstrating preparation, syndrome extraction, and logical protection for oscillator-encoded qubits [6, 7]. More recently, an integrated photonic platform has produced optical GKP states on-chip with multiple resolvable peaks and negative-Wigner structure, underscoring growing practical momentum [8].
Lattices from the geometry of numbers (see [9, 10, 11]) provide a convenient way to formalize GKP codes. In this formalism, an $n$-mode GKP code is specified by a $2n$-dimensional symplectic lattice whose vectors generate the displacement stabilizers [12, 13]. Key code parameters become geometric: in particular, the code distance—the minimum displacement magnitude that causes a logical error—is proportional to the length of shortest nonzero vector in the symplectic lattice. This lattice viewpoint also interfaces directly with channel coding, symplectically integral lattices in $2n$-dimensional phase space underlie the best-known achievable rates for the Gaussian displacement channel, motivating lattice codes spanning many oscillators [14, 15, 16, 17].
On the theory side, it is natural to call a family of $n$-mode codes good when the distance of the codes in the family grows asymptotically at least as $\Omega(\sqrt{n})$, reflecting typical high-dimensional lattice behavior where shortest-vector lengths scale as $\Omega(\sqrt{n})$ (for lattices with determinant $1$ and dimension $n$), while having asymptotically non-vanishing rate (the ratio of encoded logical qubits to physical modes is bounded away from zero) (see [18] and the references therein). Based on a non-constructive existence proof of symplectic lattices with shortest vector length $\Omega(\sqrt{n})$ by Buser and Sarnak [19], Harrington and Preskill [14] established the existence of families of good GKP codes. Complementary bounds based on continuous-variable MacWilliams identities indicate that, in low dimensions, the strongest GKP distances come from canonical ”best-packing” lattices— $E_{8}$ (8D) and the Leech lattice (24D)—widely known for maximizing minimum distance in those dimensions [20].
Constructively, Conrad, Eisert, and Seifert [18] showed that multi-mode GKP codes can be instantiated from NTRU-type lattices. NTRU is a popular lattice-based public-key cryptosystem introduced in the 1990s by Hoffstein et al. [21]. The construction by Conrad et al.  yields randomized GKP code families with non-vanishing rate and, with probability exponentially close to $1$, code distance scaling like $\Theta(\sqrt{n/q})$, where the probability is over the random choice of lattices and $q$ is some parameter depending on $n$ satisfying $n\leq q^{2}/0.28$. Exploiting so-called lattice trapdoors for NTRU lattices, Conrad et al. also describe decoding algorithms for their GKP codes that beat generic decoding procedure based on general algorithms solving the closest vector problem in lattices, such as Babai’s algorithm [22].
Note that due to the constraint $n\leq q^{2}/0.28$, the distance bound $\Theta(\sqrt{n/q})$ scales like $\Theta(\sqrt[4]{n})$ rather than $\Theta(\sqrt{n})$ as required for a good family of GKP codes. Based on the so called Gaussian Heuristic (GH) (see [18, 23, 24]) and experiments in low dimensional experiments, Conrad et al. conjecture that variants of their NTRU construction lead to families of GKP codes that encode $n\log(\lambda),\lambda>1,$ logical qubits using $n$ bosonic modes and achieve code distance $\Delta\geq\sqrt{\frac{n}{\lambda\pi e}}$ (Conjecture 1 and 2 in [18]). They also provide arguments showing that replacing NTRU lattices by more general lattices based on $q$-ary symmetric matrices leads to GKP codes with distance $\Delta\geq\sqrt{\frac{n}{\lambda\pi e}}$ (see the introduction and Proposition 3 in [18]).
In this paper, we present a simple randomized construction of multi-mode GKP codes from lattices as they appear in the (standard) short integer solutions (SIS) problem and its variants, the ring short integer solutions (R-SIS) problem and the module short integer solutions (M-SIS) problem. These lattices have been used to define e. g. collision-resistant hash functions and unforgeable signature schemes in cryptography (for an introduction see [25] although we do not use any crypto-relevant properties of these lattices). Following basically the (randomized) construction blueprint provided by [18] and using $\textrm{SIS},\textrm{R-SIS}$ and M-SIS lattices, we obtain good families of GKP codes that achieve code distance $\approx\sqrt{\frac{n}{\lambda\pi e}}$ for $n\log(\lambda)$ encoded logical qubits, that is, exactly the distances conjectured by Conrad et al. for variants of NTRU lattices. Our code families achieve these distances with probability close to $1$, where the probability is over a uniform random choice of a lattice from some finite class of lattices with dimension $n$, the set of symmetric $n$-dimensional SIS or M-SIS lattices to be precise. Choosing a lattice uniformly at random from these classes can be done efficiently, leading to an efficient construction of good GKP codes. The construction based on SIS lattices yields the most general class of GKP codes. It is also the easiest to analyze. We also provide constructions based on M-SIS lattices (R-SIS lattices are just a special case) for two reasons.
  1. 1.
    GKP codes based on M-SIS, and in particular R-SIS lattices, are closest to the NTRU construction in [18], hence they may shed more light onto those constructions,
  2. 2.
    more importantly, we also provide simple decoding algorithms for our GKP codes; using standard methods from lattice-based cryptography (e. g. fast Fourier transform), for M-SIS lattices the algorithms can be made to run in almost linear time, whereas for SIS based codes they require quadratic time.
We obtain good GKP codes by first constructing so called $q$-symplectic lattices with minimal distance $\approx\sqrt{\frac{qn}{\pi e}}$ (with high probability) from SIS lattices defined via symmetric matrices [^1] These lattices are also used by Conrad et al. for Proposition 3 to justify Conjecture 1 and 2. It is the interpretation of these lattices as SIS lattices that leads to our main results.. Here $q$ is a parameter that defines an appropriate finite field. Scaling these lattices by $\frac{1}{\sqrt{q}}$ we obtain symplectic lattices with distance $\approx\sqrt{\frac{n}{\pi e}}$. Further scaling these lattices by $\sqrt{\lambda}$ yields GKP codes with the desired distance and with $n\log(\lambda)$ logical qubits. Note that symplecticity of the lattices is needed to obtain GKP codes. As a by-product of our GKP code construction, we obtain a probabilistic construction of $n$-dimensional symplectic lattices with distance $\sqrt{\frac{n}{\pi e}}$. To the best of our knowledge, previously, the existence of such symplectic lattices was only proven non-constructively (see [19]) [^2] In [19], as well as in many other papers, results of this type are described in terms of $(1/\sigma_{2n})^{1/2n}$, where $\sigma_{2n}$ denotes the volume of the $2n$-dimensional unit ball. By Stirling’s approximation $(1/\sigma_{2n})^{1/2n}\approx\sqrt{\frac{n}{\pi e}}$..
In addition to the blueprint from [18], we also use standard techniques to analyze the minimal distance of random SIS lattices. These analyses (for SIS and M-SIS lattices) are our main technical contribution. For SIS lattices, the analysis of the minimal distance closely follows [26]). However, as mentioned above, to obtain GKP codes we cannot choose arbitrary SIS lattices. Instead, their defining matrices must be symmetric. It turns out, that this poses no problem, as we prove that random symmetric matrices over finite fields share an important regularity property with general matrices: for any fixed vector v, its image under a uniformly random matrix (general or symmetric) is also uniformly distributed. For M-SIS lattices, the symmetry issue also exists and can be resolved as for SIS lattices. However, for M-SIS lattices a more difficult technical problem arises since we have to deal with equations over finite rings rather than finite fields. More precisely, for M-SIS lattices we deal with the ring $\mathbb{Z}_{q}[X](X^{n}+1)$, where $\mathbb{Z}_{q}=\mathbb{Z}/(q\mathbb{Z})$. Although we present a very general analysis, it leads to non-trivial results only for specific choices of $n$ and $q$. The easiest and probably most useful case occurs when $n$ is a power of $2$ and $q=3,5\mod 16$, in which case the polynomial $X^{n}+1$ factors over $\mathbb{Z}_{q}[X]$ into two irreducible polynomials with the same degree. Note that these choices for $n,q$ are also popular parameter choices in cryptography and they correspond to useful choices for the NTRU-based constructions in [18].
Using standard tools from lattice theory one can show that GKP codes from symplectic lattices can not achieve distance better than $\sqrt{2n/\lambda},$ where $n$ as before is the number of modes and $n\log(\lambda)$ is the number of encoded logical qubits. This bound follows easily from so called transference bounds [27] and similar bounds were also shown in [12]. This bound shows that our construction achieves, up to small constants, the best possible distance. The Gaussian Heuiristic (see [18]) indicates that the distance of our codes are even best possible, at least as long as one insists on achieving the distance with probability close to $1$.
Finally, using the FPLLL library [28] for lattice reduction, we provide various numerical simulations for our constructions of GKP codes and our simple decoding procedure. These simulations not only confirm the theoretical findings, but also show that at least in small dimensions, more general parameter choices than supported by our theoretical results lead to GKP codes with distance $\geq\sqrt{\frac{n}{\lambda\pi e}}$. To give an example, for SIS lattices to yield good GKP codes with probability close to $1$, the theory requires $q\approx n^{2}$. The experiments show that we can choose $q$ much smaller than $n^{2}$ and still obtain codes with large distance. Although $q$ does not show up in our bounds for the code distance (it only influences probabilities), it does play an important role in our decoding algorithm. Here smaller values of $q$ lead to significantly better decoding properties. In fact, by choosing small $q$’s, our decoding procedure compares well to the NTRU decoding algorithm by Conrad et al.[18]. Similarly, our experiments show that the M-SIS-based construction is much less sensitive to choices for $n$ and $q$ than can be deduced from the theoretical results.
Organization
In Section 2 we provide all necessary concepts and results from lattice theory. The SIS-based construction of symplectic lattices is given in Section 3, the analysis is presented in Section 4, and the application to the construction of good GKP codes in Section 5. Here we also provide some well-known bounds on the invariants of symplectic lattices. In Section 6 we present our simple decoding algorithm. Section 7 summarizes our numerical findings for the SIS-based construction of GKP codes. In Sections 8 through 12 we show how to construct symplectic lattices from M-SIS lattices (Section 8), we analyze the construction (Section 9), apply it to the construction of GKP codes (Section 10), describe the adaptation of our decoding algorithm to the M-SIS-based construction (Section 11), and present and discuss the experimental results for the M-SIS-based construction (Section 12).
2 Basic definitions and results
Throughout this paper, vectors in some $\mathbb{R}^{d}$ are denoted by bold letters. A vector v is always understood as a row vector. Similarly, matrices will be denoted by capitalized bold letters. For a vector $\textbf{v}\in\mathbb{R}^{d}$, we denote its standard euclidean distance by $\|\textbf{v}\|$.
Lattices
A lattice ${\cal L}\subset\mathbb{R}^{d}$ is a discrete, abelian subgroup of the euclidean space $\mathbb{R}^{d}$. Every lattice ${\cal L}\subset\mathbb{R}^{d}$ has a basis consisting of $k,k\leq d,$ $\mathbb{R}$- linearly independent vector $\textbf{b}_{1},\ldots,\textbf{b}_{k}$, i.e., as a set,
${\cal L}:=\left\{\sum_{i=1}^{k}z_{i}\textbf{b}_{i}:z_{i}\in\mathbb{Z},i=1,\ldots,k\right\}.$
[…]
${\cal L}={\cal L}(\textbf{B})=\mathbb{Z}^{k}\textbf{B}.$
We call a lattice ${\cal L}={\cal L}(\textbf{B}),\textbf{B}\in\mathbb{R}^{k\times d}$ full-dimensional iff $k=d$.
Given some lattice ${\cal L}\subset\mathbb{R}^{d}$ its minimal distance or its first successive minimum $\lambda_{1}({\cal L})$ is defined as
$\lambda_{1}({\cal L}):=\min\{\|\textbf{v}\|:\textbf{v}\in{\cal L},\textbf{v}\neq\textbf{0}\}.$
[…]
$\rho({\cal L}):=\max\{\operatorname{dist}(\textbf{t},{\cal L})\mid\textbf{t}\in\operatorname{span}({\cal L})\},$
where for $\textbf{t}\in\operatorname{span}({\cal L})$
$\operatorname{dist}(\textbf{t},{\cal L}):=\min\{\|\textbf{t}-\textbf{u}\|\mid\textbf{u}\in{\cal L}\}.$
[…]
${\cal L}^{\dagger}:=\left\{\textbf{v}\in\operatorname{span}({\cal L})\mid\forall\textbf{u}\in{\cal L}:\textbf{v}\cdot\textbf{u}^{T}\in\mathbb{Z}\right\}.$
If ${\cal L}={\cal L}(\textbf{B})$ is full-dimensional, then ${\cal L}^{\dagger}={\cal L}(\textbf{B}^{-T})$.
The following celebrated theorem is due to Banszczyk. It is the strongest possible form (up to constants) of a transference theorem relating invariants of a lattice and its dual (see [27]) .
Theorem 1 (Banaszczyk).
[…]
$\lambda_{1}({\cal L}^{\dagger})\cdot\rho({\cal L})\leq\frac{n}{2}.$
Symplectic matrices and symplectic lattices
Set
$\textbf{J}_{2n}:=\left(\begin{matrix}\textbf{0}_{n}&\textbf{I}_{n}\\ -\textbf{I}_{n}&\textbf{0}_{n}\end{matrix}\right).$
where $\textbf{0}_{n}$ denotes the $n\times n$ matrix with all entries being 0 and $\textbf{I}_{n}$ denotes the $n\times n$ identity matrix. This matrix defines the standard symplectic form on $\mathbb{R}^{2n}$. Let $q\in\mathbb{N}$. A matrix $\textbf{M}\in\mathbb{Z}^{2n\times 2n}$ is called $q$-symplectic if
$\textbf{M}\cdot\textbf{J}_{2n}\cdot\textbf{M}^{T}=q\textbf{J}_{2n}.$
[…]
Lemma 2.
Let $\textbf{M}\in\mathbb{Z}^{n\times n}$ be a $q$-symplectic matrix. Then $\frac{1}{\sqrt{q}}\textbf{M}$ is symplectic. Accordingly, for a $q$-symplectic matrix M, lattice ${\cal L}(\frac{1}{\sqrt{q}}\textbf{M})$ is symplectic.
Definition 3.
[…]
${\cal L}^{\times}:=\left\{\textbf{v}\in\operatorname{span}({\cal L})\mid\forall\textbf{u}\in{\cal L}:\textbf{u}\cdot\textbf{J}_{2n}\cdot\textbf{v}^{T}\in\mathbb{Z}\right\},$
This is analogous to the canonical dual lattice, but instead of the euclidean inner product uses the symplectic inner product $\langle\textbf{u},\textbf{v}\rangle_{\text{symp}}=\textbf{u}\textbf{J}_{2n}\textbf{v}^{T}$. The following two lemmata are well-known and easy to verify.
Lemma 4.
[…]
$(\alpha{\cal L})^{\times}=\frac{1}{\alpha}{\cal L}.$ (1)
Lemma 5.
[…]
Proof.
[…]
Theorem 6.
[…]
$\rho((\alpha{\cal L})^{\times})=\frac{1}{\alpha}\rho({\cal L})\leq\frac{n}{2\alpha\lambda_{1}({\cal L})}.$
Proof.
The theorem follows from the following sequence of (in-)equalities.
$\rho((\alpha{\cal L})^{\times})=\frac{1}{\alpha}\rho({\cal L}^{\times})$ $=\frac{1}{\alpha}\rho({\cal L})$
$\leq\frac{1}{\alpha}\frac{n}{2\lambda_{1}({\cal L}^{\dagger})}$
$=\frac{1}{\alpha}\frac{n}{2\lambda_{1}({\cal L}^{\times})}$
$=\frac{1}{\alpha}\frac{n}{2\lambda_{1}({\cal L})}.$
[…]
A simple geometric lemma
By $B_{d}(r)$ denote the $d$-dimensional ball of radius $r$ centered at the origin. The volume of $B_{d}(1)$ is denoted by $\sigma_{d}$. The next lemma is well-known. Below, it plays a crucial role in the probabilistic analysis of the minimum distance of random symplectic lattices. For the sake of completeness we give a proof.
Lemma 7.
Let $d\in\mathbb{N}$ and $r$ be a positive real number. Then
$\left|\mathbb{Z}^{d}\cap B_{d}(r)\right|\leq\sigma_{d}(r+\sqrt{d}/2)^{d}.$
Proof.
Around each integer point $z$ in $B_{d}(r)$ place an open $d$-dimensional cube centered at $z$ with side-length $1$, denoted by $C(z)$. Every cube $C(z)$ has $d$-dimensional volume $1$. The cubes around different integer points do not intersect. Moreover, for all $z$ we have $C(z)\subset B_{d}(r+\sqrt{d}/2)$. Denote the $d$-dimensional volume of a (measurable) set $S$ by $V_{d}(S)$. Then from the above we deduce
$\left|\mathbb{Z}^{d}\cap B_{d}(r)\right|=V_{d}\left(\bigcup_{z\in\mathbb{Z}^{d}\cap B_{d}(r)}C(z)\right)\leq$ $V_{d}\big{(}B_{d}(r+\sqrt{d}/2)\big{)}^{d}$
$=$ $\sigma_{d}\cdot\big{(}r+\sqrt{d}/2\big{)}^{d}.$
Some useful bounds
The probabilistic analysis of crucial parameters of random symplectic lattices below also uses the following well known bounds on factorials and volumes of high-dimensional balls.
Lemma 8.
As before denote by $\sigma_{d}$ the volume of the unit ball $B_{d}(1)$ in $\mathbb{R}^{d}$. Then
  • (1)
    $\sqrt{2\pi d}\left(\frac{d}{e}\right)^{d}e^{\frac{1}{12d+1}}\leq d!\leq\sqrt{2\pi d}\left(\frac{d}{e}\right)^{d}e^{\frac{1}{12d}}$
  • (2)
    $\sqrt{2\pi d}^{1/(2d)}\sqrt{\frac{d}{\pi e}}\leq\left(\frac{1}{\sigma_{2d}}\right)^{1/(2d)}\leq\sqrt{2\pi d}^{1/(2d)}\sqrt{\frac{d}{\pi e}}e^{\frac{1}{12d^{2}}}$
The first statement of the lemma is an explicit version of Stirling’s formula due to Robbins [29]. The second one follows from the first statement and $\sigma_{2d}=\frac{\pi^{d}}{d!}$. The second statement of the lemma shows that up to low order terms $(1/\sigma_{2d})^{1/2d}$ and $\sqrt{d/\pi e}$ coincide. More precisely,
$\left(\frac{1}{\sigma_{2d}}\right)^{1/(2d)}=\left(1+o\bigl{(}\frac{\ln(d)}{d}\bigr{)}\right)\sqrt{\frac{d}{\pi e}}.$ (2)
3 Constructing symplectic matrices from SIS lattices
[…]
So called modular lattices play an important role in modern lattice-based cryptography. In the most general case, such a lattice is defined via a matrix $\textbf{A}\in\mathbb{Z}^{n\times m},m>n,$ and an integer $q$. Then one sets
${\cal L}^{\perp}(\textbf{A}):=\{\textbf{v}\in\mathbb{Z}^{m}:\textbf{A}\cdot\textbf{v}=0\mod q\}.$
One easily checks, that this indeed defines a lattice. These lattices appear in cryptography in the short integer solutions problem (SIS). Mostly for efficiency reasons, nowadays cryptographers prefer modular ring lattices. As it turns out, these are also useful in the explicit construction of symplectic lattices and GKP codes. NTRU lattices as used by Conrad et al [18], in many ways are just special modular ring lattices. We define these lattices later.
Let $\textbf{H}\in\mathbb{Z}^{n\times n}$ be a symmetric matrix, i.e. $\textbf{H}=\textbf{H}^{T}$, and let $q\in\mathbb{N}$. Consider the $2n\times 2n$ matrix
$M_{\textbf{H}}=\left(\begin{matrix}\textbf{I}_{n}&\textbf{H}\\ \textbf{0}_{n}&q\textbf{I}_{n}\end{matrix}\right).$ (3)
A simple calculation shows that $M_{\textbf{H}}$ is $q$-symplectic. The rows of $M_{\textbf{H}}$ generate a $2n$-dimensional lattice ${\cal L}_{\textbf{H}}={\cal L}(M_{\textbf{H}})$, i.e. a $q$-symplectic lattice.
Next, define
$A_{\textbf{H}}:=\left(\textbf{H}\mid-\textbf{I}_{n}\right)\in\mathbb{Z}^{n\times 2n}$
and consider lattice ${\cal L}^{\perp}(A(\textbf{H}))$. It has a basis given by the rows of matrix $M_{\textbf{H}}$ defined above, i.e.
${\cal L}^{\perp}(\textbf{A}(\textbf{H}))={\cal L}(M_{\textbf{H}}).$ (4)
Simplifying notation, we set
${\cal L}^{\perp}(\textbf{H}):={\cal L}^{\perp}(\textbf{A}(\textbf{H})).$ (5)
As noted above, if H is symmetric, then ${\cal L}^{\perp}(\textbf{H})$ is $q$-symplectic and $1/\sqrt{q}{\cal L}^{\perp}(\textbf{H})$ is symplectic. We call lattices of the form $1/\sqrt{q}{\cal L}^{\perp}(\textbf{H}),$ H symmetric, symplectic SIS lattices.
Symplectic SIS lattices are also used by Conrad et al. (see Proposition 3 in [18]). However, they did not interpret them as SIS lattices.
4 Analysis of minimal distance
In this section we prove that with probability close to $1$ lattices ${\cal L}^{\perp}(\textbf{H})$ have minimal distance $\approx\sqrt{qn/\pi e}$. Here the probability is with respect to the uniform distribution over symmetric matrices H in $\mathbb{Z}_{q}^{n\times n},n\in\mathbb{N}$.
Definition 9.
For $n\in\mathbb{N}, qq$ a prime power, denote the set of symmetric matrices in $\mathbb{F}_{q}^{n\times n}$ by $\operatorname{Sym}(q,n)$.
To generate an element of $\operatorname{Sym}(q,n)$ uniformly at random proceed in two steps
  1. 1.
    choose $n+1\choose 2$} elements $h_{ij},1\leq i,j\leq n,j\geq i.$, uniformly at random from $\mathbb{F}_{q}$,
  2. 2.
    for $1\leq i\leq n,1\leq j<i$, set $h_{ij}=h_{ji}$
The analysis of the minimal distance proceeds in two steps and follows standard arguments in lattice-based cryptography (see for example [26]), but taking into account that we work with symmetric matrices. In the first step of the analysis, we determine for fixed $\textbf{z}\in\mathbb{Z}^{2kn},$
$\Pr[\textbf{z}\in{\cal L}^{\perp}(\textbf{H}):\textbf{H}\leftarrow\operatorname{Sym}(q,n)].$ (6)
Then we use an upper bound on short elements in $\mathbb{Z}^{2n}$ (see Lemma 7) and the union bound to obtain the main result.
4.1 Analyzing the probability
First, we need a general lemma on solutions to random symmetric linear equations. For $q$ a prime power, denote by $\mathbb{F}_{q}$ the field with $q$ elements.
Lemma 10.
Let $\textbf{z}_{i}\in\mathbb{F}_{q}^{2n},i=1,2,\textbf{z}=(\textbf{z}_{1},\textbf{z}_{2})\neq\textbf{0}$. Then
$\Pr[\textbf{H}\textbf{z}_{1}=\textbf{z}_{2};\textbf{H}\leftarrow\operatorname{Sym}(q,n)]=\frac{1}{q^{n}}.$
Note that this is the same bound that we obtain by choosing H uniformly at random from $\mathbb{Z}_{q}^{n\times n}$.
Proof.
If $\textbf{z}_{1}=\textbf{0}$, then $\textbf{z}_{2}=\textbf{0}\mod q,$ which is impossible since $\textbf{z}\neq\textbf{0}$ by assumption. For any permutation matrix P and symmetric H, matrix $\textbf{P}^{T}\textbf{H}\textbf{P}$ is symmetric. Therefore, by permuting the rows and columns of H and $\textbf{z}_{1},\textbf{z}_{2}$, we may without loss of generality assume that the last $m$ entries of $\textbf{z}_{1}$ are non-zero and the remaining entries are zero. Next, write a symmetric matrix $\textbf{H}\in\mathbb{F}_{q}^{n\times n}$ as
$\textbf{H}=\left(\begin{matrix}\textbf{H}_{11}&\textbf{H}_{12}\\ \textbf{H}_{21}&\textbf{H}_{22}\end{matrix}\right),$
where
  • $\textbf{H}_{11}\in\mathbb{Z}_{q}^{n-m\times n-m},\textbf{H}_{22}\in\mathbb{F}_{q}^{m\times m}$ are symmetric
  • $\textbf{H}_{21}=\textbf{H}_{12}^{T}\in\mathbb{F}_{q}^{m\times n-m}$.
Assume that all elements in H are fixed, except the elements on the diagonal of $\textbf{H}_{22}$ and the last column of $\textbf{H}_{12}$ (which is also the last row of $\textbf{H}_{21}$.) Hence, overall $n$ elements are not fixed. Then, there is a unique choice in $\mathbb{F}_{q}^{n}$ for the elements not fixed such that $\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod q$. This proves the lemma. ∎
Lemma 11.
Let $q$ be prime and $\textbf{z}\in\mathbb{Z}^{2n},\textbf{z}\neq\textbf{0}\mod q$. Then
$\Pr[\textbf{z}\in{\cal L}^{\perp}(\textbf{H});\textbf{H}\leftarrow\operatorname{Sym}(q,n)]=\frac{1}{q^{n}}.$
Proof.
Write $\textbf{z}=(\textbf{z}_{1},\textbf{z}_{2}),\textbf{z}_{1},\textbf{z}_{2}\in\mathbb{Z}^{n}$. Then
$\textbf{z}\in{\cal L}^{\perp}(\textbf{H})\Leftrightarrow\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod q.$
The lemma follows from the previous lemma. ∎
4.2 Bounding the minimal distance
Theorem 12.
Let $n,q,r\in\mathbb{N}$, where $q$ is prime and $r\sqrt{q}<q/2$. Then
$\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H})<r\sqrt{q};\textbf{H}\leftarrow\operatorname{Sym}(q,n)]\leq\sigma_{2n}\big{(}r+\sqrt{\frac{n}{2q}}\big{)}^{2n}.$ (7)
Proof.
To simplify notation, define event
$E:=\{\textbf{H}\in\mathbb{Z}_{q}^{n\times n}:\lambda_{1}(\textbf{H})<r\sqrt{q}\},$
and denote by $\bar{E}$ its complimentary event. Hence, we need to prove that
$\Pr[E(\textbf{H});\textbf{H}\leftarrow\operatorname{Sym}(q,n)]\leq\sigma_{2n}\big{(}r+\sqrt{\frac{n}{2q}}\big{)}^{2n}.$
By the union bound
$\begin{split}\Pr[\textbf{H}\in E;\textbf{H}\leftarrow\operatorname{Sym}(q,n)]&=\Pr[\exists\textbf{v}\in B_{2n}(r\sqrt{q})\cap{\cal L}^{\perp}(\textbf{H})\setminus\{\textbf{0}\};\textbf{H}\leftarrow\operatorname{Sym}(q,n)]\\ &\leq\sum_{\textbf{v}\in B_{2n}(r\sqrt{q})\cap\mathbb{Z}^{2n}\setminus\{\textbf{0}\}}\Pr[\textbf{v}\in{\cal L}^{\perp}(\textbf{H});\textbf{H}\leftarrow\operatorname{Sym}(q,n)].\end{split}$
By Lemma 7
$|B_{2n}(r\sqrt{q})\cap\mathbb{Z}^{2n}\setminus\{\textbf{0}\}|\leq\sigma_{2n}\big{(}r\sqrt{q}+\frac{\sqrt{2n}}{2}\big{)}^{2n}=\sigma_{2n}\big{(}r+\sqrt{\frac{n}{2q}}\big{)}^{2n}q^{n}.$
Since $r\sqrt{q}<q/2$, elements in $B_{2n}(r\sqrt{q})\cap\mathbb{Z}^{2n}\setminus\{\textbf{0}\}$ are $\neq\textbf{0}\mod q$. Hence, combining the upper bound on $\Pr[\textbf{H}\in E;\textbf{H}\leftarrow\operatorname{Sym}(q,n)]$ with Lemma 11 proves the theorem. ∎
Corollary 13.
Let $n,q\in\mathbb{N},r\in\mathbb{R}^{>0}$, where $q$ is prime and $r<\sqrt{q}/2$. Choosing H uniformly at random from $\operatorname{Sym}(q,n)$ and setting
${\cal L}:=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H})$
[…]
$p:=\sigma_{2n}\big{(}r+\sqrt{\frac{n}{2q}}\big{)}^{2n}$
[…]
Proof.
By Lemma 2, ${\cal L}=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H})$ is symplectic. Its minimal distance $\lambda_{1}({\cal L})$ is given by $\frac{1}{\sqrt{q}}\lambda_{1}({\cal L}^{\perp}(\textbf{H})).$ The corollary follows from Theorem 12 . ∎
5 Main results
In this section we discuss several parameter choices in Theorem 12. Some of these choices lead to a constructive versions of a theorem by Buser and Sarnak [19] on the existence of symplectic lattices with bounded minmal distance, albeit with slightly worse constants than in [19]. Other choices lead to variants of a conjecture by Conrad, Eisert and Seifert [18] on the existence and construction of good GKP codes. In both cases, different parameter choices lead to different combinations of success probability and minimal lattice or code distance.
5.1 Symplectic lattices with large minimal distance
Theorem 14 (Existence of symplectic lattices with large minimal distance).
Choose H uniformly at random from $\operatorname{Sym}(q,n)$ and set
${\cal L}:=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H}).$
  1. 1.
    For any $\epsilon>0$ and $q\geq\frac{2n\cdot\sigma_{2n}^{1/n}}{\epsilon^{2}}$, with non-zero probability
    $\lambda_{1}({\cal L})\geq(1-\epsilon)(1/\sigma_{2n})^{1/2n},$
    where $\sigma_{2n}$ denotes the volume of the unit ball in $\mathbb{R}^{2n}$.
  2. 2.
    For $q\geq\frac{\pi en^{2}}{2}$, with probability at least $1-\frac{e^{2}}{\sqrt{2\pi n}}=1-\Omega(1/\sqrt{n})$
    $\lambda_{1}({\cal L})\geq\sqrt{\frac{n}{\pi e}},$
  3. 3.
    For $q\geq 2\pi en^{3/2}$, with probability at least $1-e^{-n^{1/4}}$
    $\lambda_{1}({\cal L})\geq\left(1-\frac{1}{n^{3/4}}\right)\sqrt{\frac{n}{\pi e}}.$
Proof.
For the first part of the theorem, use Corollary 13 with
$p:=(1-\epsilon/2)^{2n}$
[…]
$\lambda_{1}({\cal L})\geq(1-\epsilon)(1/\sigma_{2n})^{1/2n}.$
For the second part of the theorem, set
$p:=\frac{e^{2}}{\sqrt{2\pi n}}$
and apply Corollary 13. This shows that with probability at least 1p1-p the lattice ${\cal L}={\cal L}^{\perp}(\textbf{H})$ has distance
$\lambda_{1}({\cal L})\geq(p/\sigma_{2n})^{1/2n}-\sqrt{n/2q}.$
By the choice of $q$
$\sqrt{n/2q}\leq\frac{1}{n}\sqrt{\frac{n}{\pi e}}.$
Using Lemma 8 (2) and by our choice of $p$,
$(p/\sigma_{2n})^{1/2n}\geq e^{1/n}\sqrt{\frac{n}{\pi e}}\geq\left(1+\frac{1}{n}\right)\sqrt{\frac{n}{\pi e}}.$
The third part of the theorem follows as the previous two parts, by choosing
$p:=e^{-n^{1/4}}$
and using
$(p/\sigma_{2n})^{1/2n}\geq\left(1-\frac{1}{2n^{3/4}}\right)\sqrt{\frac{n}{\pi e}}\quad\text{and $\sqrt{\frac{n}{2q}}\leq\frac{1}{2n^{3/4}}\sqrt{\frac{n}{\pi e}}$}.$
Discussion
  • Buser and Sarnak [19] prove the existence of $2n$-dimensional lattices with first successive minimum
    $\lambda_{1}({\cal L})\geq(2/\sigma_{2n})^{1/2n}\approx(1+1/2n)(1/\sigma_{2n})^{1/2n}.$
    Since we can take $\epsilon>0$ arbitrary, the bound by Buser and Sarnak is better than the bound in the first part of Theorem 14 by a factor of $\approx 1+1/2n$. The other two parts show that from symplectic SIS lattices we can get, with high probability, symplectic lattices with distance close to $\sqrt{n/\pi e}$, which differs from the Buser, Sarnak bound only by small factors, approaching $1$ with $n\rightarrow\infty$.
  • The lower bounds on $q$ in Theorem 14 are required to guarantee probabilities close to $1$. Ultimately, they can be traced back to the additive terms of $\sqrt{d}/2$ in Lemma 7. Many researchers believe that this additional term of $\sqrt{d}/2$ is not necessary, or can be replaced by something much smaller. If we can ignore it, then the lower bounds on $q$ can be reduced to roughly $n$ (going back to the requirement $r\sqrt{q}/2<q$ in Theorem 12). The experiments and simulations that we present in Section 7 confirm that we can choose $q$ much smaller than indicated by Theorem 14 and still obtain symplectic lattices and GKP with large minimal distance.
5.2 GKP codes with large distance
[…]
$\forall\textbf{u},\textbf{v}\in{\cal L}:\textbf{u}\textbf{J}_{2n}\textbf{v}^{T}\in\mathbb{Z}.$
[…]
$\Delta({\cal C}):=\min\{\|\textbf{u}\|\mid\textbf{u}\in{\cal L}^{\times}\setminus{\cal L}\},$
[…]
$\Delta({\cal C})\geq\lambda_{1}({\cal L}^{\times}).$
[…]
$\Delta({\cal C})=\lambda_{1}((\sqrt{\lambda}{\cal L})^{\times})=\frac{1}{\sqrt{\lambda}}\lambda_{1}({\cal L}).$ (8)
(see [12] and Lemma 4). Equation (8) together with Theorem 14, part 2 and 3, yields the following theorem.
Theorem 15 (Good GKP codes from SIS lattices).
A GKP code $\cal C$} with stabilizer group ${\cal L}=\sqrt{\lambda/q}{\cal L}^{\perp}(\textbf{H})$, where ${\cal L}^{\perp}(\textbf{H})$ is as in (5) has $\lambda^{n}$ logical dimensions. Over the choice of $\textbf{H}\leftarrow\operatorname{Sym}(q,n)$, the distance $\Delta({\cal C})$ of $\cal C$} satisfies the following.
  1. 1.
    For $q\geq\frac{\pi en^{2}}{2}$, with probability at least $1-\frac{e^{2}}{\sqrt{2\pi n}}=1-\Omega(1/\sqrt{n})$
    $\Delta({\cal C})\geq\sqrt{\frac{n}{\lambda\pi e}},$
  2. 2.
    For $q\geq 2\pi en^{3/2}$, with probability at least $1-e^{-n^{1/4}}$
    $\Delta({\cal C})\geq\left(1-\frac{1}{n^{3/4}}\right)\sqrt{\frac{n}{\lambda\pi e}}.$
Discussion Conrad et al.  [18] conjecture that certain GKP code constructions based on NTRU lattices from cryptography yield (with high probability) GKP codes with distance $\sqrt{n/\lambda\pi e}$. Part one of the above theorem shows that GKP codes with distance $\geq\sqrt{n/\lambda\pi e}$ can be constructed from SIS lattices. Part two shows that, by going slightly below $\sqrt{n/\lambda\pi e}$, we can achieve distance roughly $\sqrt{n/\lambda\pi e}$ from SIS lattices with probability exponentially close to $1$.
5.3 Upper bounds for minimal distance and covering radius and
In this section, we prove some upper bounds on the minimal distance of symlectic SIS lattices and GKP codes obtained from such lattices, as well as on the covering radius of symplectic SIS lattices. All bounds are either well-known or can be proven easily using techniques known from general SIS lattices. For the sake of completeness we include proofs.
Lemma 16.
[…]
  1. 1.
    $\lambda_{1}({\cal L})\leq\sqrt{2n}$
  2. 2.
    $\lambda_{1}((\sqrt{\lambda}{\cal L})^{\times})\leq\sqrt{\frac{2n}{\lambda}}$.
Proof.
[…]
$\lambda_{1}({\cal L})\leq\sqrt{2n}\det({\cal L})^{1/2n}=\sqrt{2n}.$
This proves the first part of the lemma. The second follows from part 1 and Lemma 4. ∎
The next corollary follows from the previous lemma and (8).
Corollary 17.
[…]
$\Delta({\cal C})\leq\sqrt{\frac{2n}{\lambda}}.$
Theorem 18.
Let $\lambda\in\mathbb{N}$. Choose H uniformly at random from $\operatorname{Sym}(q,n)$ and set
${\cal L}:=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H}).$
  1. 1.
    For $q\geq\frac{\pi en^{2}}{2}$, with probability at least $1-\frac{e^{2}}{\sqrt{2\pi n}}=1-\Omega(1/\sqrt{n})$
    $\rho((\sqrt{\lambda}{\cal L})^{\times})\leq\frac{1}{2}\sqrt{\frac{\pi en}{\lambda}}.$
  2. 2.
    For $q\geq 2\pi en^{3/2}$, with probability at least $1-e^{-n^{1/4}}$
    $\rho((\sqrt{\lambda}{\cal L})^{\times})\leq\left(1-\frac{1}{n^{3/4}}\right)^{-1}\frac{1}{2}\sqrt{\frac{\pi en}{\lambda}}.$
Proof.
[…]
$\rho((\sqrt{\lambda}{\cal L})^{\times})=\frac{1}{\sqrt{\lambda}}\rho({\cal L}^{\times})\leq\frac{n}{2\sqrt{\lambda}\lambda_{1}({\cal L})}.$
Hence the first part of the theorem follows from part two of Theorem 14. ∎
[…]
Together, Theorem 14, Lemma 16 and Theorem 18, show that for a random symplectic SIS lattice important lattice parameters like the minimal distance and the covering radius differ from each other and (the general upper bound for $\lambda_{1}({\cal L})$ of) $\sqrt{2n}$ by constants (and with high probability). Theorem 15 and Corollary 17, on the other hand, show that the distance of a GKP constructed from random symplectic SIS lattices differs from the general upper bound $\sqrt{\frac{2n}{\lambda}}$ only by constant factors (with high probability).
6 Bounded distance decoding
[…]
$\operatorname{dist}(\textbf{t},{\cal L}):=\min\{\|\textbf{t}-\textbf{u}\|\mid\textbf{u}\in{\cal L}\}.$
Definition 19 (Problem $\textrm{BDD}_{R}$).
Given an basis B for lattice ${\cal L}={\cal L}(\textbf{B})$, a distance bound $R\in\mathbb{R}^{\geq 0}$ and a point $\textbf{t}\in\operatorname{span}({\cal L})$, such that $\operatorname{dist}(\textbf{t},{\cal L})\leq R$, find a lattice point $\textbf{v}\in{\cal L}$ with $\|\textbf{t}-\textbf{v}\|\leq R$. We denote this by $\textrm{BDD}({\cal L},\textbf{t},R)$.
Often, $R$ is expressed as a multiple of $\lambda_{1}({\cal L})$ or as a multiple of its covering radius $\rho({\cal L})$. Note, that if $R<\lambda_{1}({\cal L})/2$ then a solution to $\textrm{BDD}({\cal L},\textbf{t},R)$ is unique.
To solve the BDD problem for lattices $(\sqrt{\lambda}{\cal L})^{\times}=\frac{1}{\sqrt{\lambda}}{\cal L},$ where ${\cal L}=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H})$ is symplectic, we propose the following simple algorithm BDD. The input to the algorithm is a symmetric matrix $\textbf{H}\in\mathbb{Z}^{n\times n}$, a prime $q$, an integer $\lambda$, together defining the lattice $\frac{1}{\sqrt{\lambda q}}{\cal L}^{\perp}(\textbf{H})$, and a target vector $v=(v_{1},\ldots,v_{n},v_{n+1},\ldots,v_{2n})$. It outputs an integer vector $c=(c_{1},\ldots,c_{n},c_{n+1},\ldots,c_{2n})$. Implicitly, the algorithm uses the basis $\frac{1}{\sqrt{\lambda q}}M(\textbf{H})$ for lattice $\frac{1}{\sqrt{\lambda q}}{\cal L}^{\perp}(\textbf{H})$ (see (4) and (5)).
$\textsc{Bdd}_{\text{triv}}$ on input $\textbf{H}\in\mathbb{Z}^{n\times n},q,\lambda\in\mathbb{N}$ and $\textbf{v}\in\mathbb{Z}^{2n}$: 1. For $i=1,\ldots,n$, set $c_{i}:=\lfloor\sqrt{\lambda q}v_{i}\rceil$. 2. Set $\textbf{c}^{\prime}:=(c_{1},\ldots,c_{n}),\textbf{w}=\textbf{v}^{\prime}-\textbf{c}^{\prime}\textbf{H}$, where $\textbf{v}^{\prime}:=(v_{n+1},\ldots,v_{2n})$, set $\textbf{w}=(w_{1},\ldots,w_{n})$. 3. For $i=n+1,\ldots,2n$, set $c_{i}:=\lfloor\sqrt{\lambda/q}w_{n-i}\rceil$. 4. Output $\textbf{c}=(c_{1},\ldots,c_{2n})$.
The next theorem follows directly from the algorithm.
Theorem 20.
[…]
If we compare the bound $\frac{1}{2\sqrt{\lambda q}}$ from Theorem 20 with
  • the lower bounds on the minimal distance and covering radius of lattice ${\cal L}=\frac{1}{\sqrt{\lambda q}}{\cal L}^{\perp}(\textbf{H})$ from Theorem 14 (part three) and from Theorem 18 (part three), respectively
  • the lower bound on the minimal distance of GKP code $\cal C$} with stabilizer $\sqrt{\frac{\lambda}{q}}{\cal L}^{\perp}(\textbf{H})$ from Theorem 15 (part two)
we see that algorithm $\textsc{Bdd}_{\text{triv}}$
  1. 1.
    computes the closest lattice vector to target vector t if
    • $\operatorname{dist}(\textbf{t},{\cal L})\leq\gamma\lambda_{1}({\cal L})$ for $\gamma\leq\frac{(1-1/n^{3/4})^{-1}}{2\sqrt{2}n^{5/4}}$,
    • $\operatorname{dist}(\textbf{t},{\cal L})\leq\gamma\rho({\cal L})$ for $\gamma\leq\frac{(1-1/n^{3/4})^{-1}}{\sqrt{2}\pi en^{5/4}}$,
  2. 2.
    computes the closest code word to target vector t if $\operatorname{dist}(\textbf{t},{\cal C})\leq\gamma\Delta({\cal C})$ for $\gamma\leq\frac{(1-1/n^{3/4})^{-1}}{2\sqrt{2}n^{5/4}}$,
in all cases with probability exponentially close to $1$ over the random choice of H.
7 Simulation of main results
In this section we show some simulation results corresponding to the main results in the above sections. The simulation results shows the correctness of the main results from another perspective.
7.1 Environment setting
[…]
7.2 Distance of GKP code from SIS lattices
[…]
(a) fix $q=256$
(b) fix $n=7$
Figure 1: Code distance of GKP code generated from SIS lattice, (a) with $q=256$ fixed, vary $n$ as horizontal axis; (b) with $n=7$ fixed, vary $q$ as horizontal axis.
Fig. 1 shows two cases of code distance varying with $n$ and $q$. In the two graphs, the maximum, median and minimum of code distance of 1000 GKP code samples are plotted. The numbers below each data point in the plot show the rate of samples with code distance $\Delta>\sqrt{n/2\pi e}$. Either in Fig. 1 (a) or Fig. 1 (b), more than 90 percent samples have code distance above the bound $\sqrt{n/2\pi e}$, which correspond to Theorem 15. It is worth nothing that this high above bound rate is for all combinations of $(n,q)$ in the simulation, which gives the conjecture that with the construction given by Theorem 15, code distance have high probability (close to 1) to be more than the bound $\sqrt{n/\lambda\pi e}$ for all $(n,q)$ combination, i.e. also for small $q\leq\pi en^{2}/2$ or $q\leq 2\pi en^{3/2}$. One can also see from Fig. 1 that Corollary 17 and Theorem 18 are true in the simulation.
7.3 Trivial decoder for BDD
In this section we show the decoding performance of the trivial decoder $\textsc{Bdd}_{\text{triv}}$, as well as comparing it with NTRU decoder mentioned in [12] and Babai decoder using Babai’s algorithm [22].
[…]
(a) trivial decoder, $n=11$
(b) trivial decoder, $q=4$
(c) decoder comparison, $n=7$, $q=4$
(d) decoder comparison, $n=7$, $q=8$
Figure 2: Decoding under Gaussian displacement noise. (a) Trivial decoder at $n=11$ for $q\in\{4,5,7,8,16\}$. (b) Trivial decoder at $q=4$ for $n\in\{7,11,17,23,31\}$. (c–d) NTRU decoder, Babai decoder, and trivial decoder ($\textsc{Bdd}_{\text{triv}}$) at $n=7$ with $q=4$ and $q=8$. Curves show error rate $p_{\mathrm{err}}$ versus physical standard deviation $\sigma$. For each $(n,q)$, we select the code with the largest distance from 100 random candidates and estimate $p_{\mathrm{err}}$ using $10^{4}$ error samples per $\sigma$. For (c–d), codes are generated from NTRU lattices to enable the NTRU decoder.
Fig. 2 (a) and (b) shows the decoding performance of $\textsc{Bdd}_{\text{triv}}$. From Fig. 2 (a) we can see that the decoding performance for trivial decoder gets worse as $q$ increase, which corresponds with Theorem 20, since the trivial decoder only computes the closest vector correctly when $\operatorname{dist}(\textbf{v},{\cal L}^{\perp})<\frac{1}{2\sqrt{\lambda q}}$. Notice that $\operatorname{dist}(\textbf{v},{\cal L}^{\perp})$ represents the amplitude of the real error. Meanwhile in our simulation, errors are generated by Gaussian error model. Therefore as $q$ increase, $\frac{1}{2\sqrt{\lambda q}}$ becomes smaller. There’s less probability to have $\operatorname{dist}(\textbf{v},{\cal L}^{\perp})<\frac{1}{2\sqrt{\lambda q}}$ for an random Gaussian error, thus the decoding performance drops.
From Fig. 2 (b), we can see that the decoding performance for the trivial decoder gets worse as $n$ increases. This also corresponds with Theorem 20, since $\gamma$ decreases as $n$ increases. Therefore, the decoding performance for the trivial decoder gets worse for similar reasons as mentioned above.
Fig. 2 (c) and (d) show two cases of the decoding performance of $\textsc{Bdd}_{\text{triv}}$, NTRU decoder and Babai decoder. Based on the simulation results, the trivial decoder performs better than both NTRU and Babai decoder when $q$ is small ($q\leq 4$), and performs worse than NTRU but still better than Babai decoder when $q$ is large. Considering that there’s no need for picking $q$ large for GKP codes (Since the distance bound $\sqrt{n/2\pi e}$ is irrelevant with $q$) and that the trivial decoder requires less computing resources, the trivial decoder can be considered as a good option for decoding GKP codes generated from lattices.
8 Symplectic lattices from R-SIS and M-SIS lattices
In this section we construct symplectic lattices from so-called ring and module SIS lattices, R-SIS and M-SIS lattices, respectively. Both are special cases of modular lattices. Furthermore, R-SIS lattices are a subclass of M-SIS lattices. We consider R-SIS lattices separately to increase readability.
8.1 Basic definitions and results
To define ring and module lattices, fix some integer polynomial $\Phi(X)$ and an integer $q$. These define rings
$R:=\mathbb{Z}[X]/(\Phi(X))\quad\text{and}\quad R_{q}:=\mathbb{Z}_{q}[X]/(\Phi(X)).$
Let $\textbf{A}\in R_{q}^{k\times m}$, for integers $k,m$. Similar, to the definition for modular lattices, set
${\cal L}^{\perp}(\textbf{A}):=\{\textbf{v}\in R^{m}:\textbf{A}\cdot\textbf{v}=0\mod qR\}.$
[…]
Using matrix representations of multiplications by ring elements, one can see that module lattices as just described are special cases of modular lattices. These matrix representations are crucial for the construction of symplectic lattices and good GKP codes.
Matrix representations
We consider $\mathbb{Z}[X]$ as a finite dimensional $\mathbb{Q}$ vector space. Throughout this paper we fix the basis $1,X,\ldots,X^{n-1}$. Let $h(X)=\sum_{i=0}^{n-1}h_{i}X^{i}$ be an element in $R:=\mathbb{Z}[X]/(\Phi(X))$. Consider the linear map
$\begin{array}[]{ll}\rho(\textbf{h}):R&\longrightarrow R\\ b&\longmapsto b\cdot h.\end{array}$ (9)
As a linear map it has a matrix representation with respect to the basis $1,\ldots,X^{n-1}$, which we also denote by $\rho(\textbf{h})$.
The most important rings considered in this paper are rings $R,R_{q}$ for the polynomials $\phi(X)=X^{n}\pm 1$. Let $h=h(X)=\sum_{i=0}^{n-1}h_{i}X^{i}$ be an element of $R$. For $\Phi(X)=X^{m}-1$ the matrix representation of $\rho(\textbf{h})$ is given by
$\rho(\textbf{h})=\left(\begin{matrix}h_{0}&h_{1}&h_{2}&\ldots&h_{n-1}\\ h_{n-1}&h_{0}&h_{1}&\ldots&h_{n-2}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ h_{2}&h_{3}&h_{4}&\ldots&h_{1}\\ h_{1}&h_{2}&h_{3}&\ldots&h_{0}\end{matrix}\right).$
For $\Phi(X)=X^{n}+1$ the matrix representation is given by
$\rho(\textbf{h})=\left(\begin{matrix}h_{0}&h_{1}&h_{2}&\ldots&h_{n-1}\\ -h_{n-1}&h_{0}&h_{1}&\ldots&h_{n-2}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ -h_{2}&-h_{3}&-h_{4}&\ldots&h_{1}\\ -h_{1}&-h_{2}&-h_{3}&\ldots&h_{0}\end{matrix}\right).$ (10)
Factorization of cyclotomic polynomials
The structure of rings $R=\mathbb{Z}[X]/(\Phi(X))$ and $R_{q}=\mathbb{Z}_{q}[X]/(\Phi(X))$ is determined by the factorization of the polynomial $\Phi(X)$ over $\mathbb{Q}$ modulo $q$. More precisely, if
$\Phi(X)=\prod_{i=1}^{m}p_{j}(X)\mod q$
for irreducible polynomials over $\mathbb{F}_{q}$ with $\deg(p_{j})=n_{j}$, then
$R_{q}\cong\bigotimes\mathbb{F}_{q}/(p_{j}(X))\cong\bigotimes_{i=1}^{m}\mathbb{F}_{q^{n_{j}}}.$ (11)
In the sequel, we only consider $\Phi(X)=X^{n}\pm 1$. Following [31], by $Q_{d}$ denote the $d$-th cyclotomic polynomial over some field $K$, that is,
$Q_{d}(X):=\prod_{\begin{subarray}{c}s=1\\ \gcd(s,d)=1\end{subarray}}^{d}(X-\zeta^{s}),$
where $\zeta$ denotes a primitive $d$-th root of unity over $K$. In our cases, $K=\mathbb{Q}$ or $K=\mathbb{F}_{q}$, where $q$ is prime and $\mathbb{F}_{q}$ denotes the field with $q$ elements. In any case, the degree $\deg(Q_{d}(X))$ is $\phi(d),$ where $\phi(\cdot)$ denotes Euler’s totient function. Then, over $\mathbb{Q}$ and over $\mathbb{F}_{q}$ and assuming $\gcd(n,q)=1$,
$X^{n}-1=\prod_{d\mid n}Q_{d}(X)\quad\text{and}\quad X^{n}+1=\prod_{\begin{subarray}{c}d\mid 2n\\ d\nmid n\end{subarray}}Q_{d}(X).$ (12)
In both cases, the coefficients of the polynomials $Q_{d}(X)$ are in $\mathbb{Q}$ and $\mathbb{F}_{q}$, respectively. Whereas in the case $K=\mathbb{Q}$ the polynomials $Q_{d}(X)$ are irreducible, this is, in general, not true for $\mathbb{F}_{q}$. Instead (see e.g. Theorem 2.47 in [31])
Theorem 21.
Let $K=\mathbb{F}_{q}$ and assume that $\gcd(d,q)=1$. Then $Q_{d}(X)$ factors into $\phi(d)/\ell$ distinct irreducible polynomial of degree $\ell$, where $\ell$ is the order of $q$ modulo $d$, i.e. $\ell$ is the smallest integer with $q^{\ell}=1\mod d$.
In applying this theorem below we use the Carmichael function $\lambda(\cdot)$, where for $n\in\mathbb{N}$, $\lambda(n)$ is the smallest integer $m$ such that
$\forall a\in\mathbb{Z}_{n}^{*}:a^{m}=1\mod n.$
Equivalently, $\lambda(n)$ is the largest integer $m$ such $\mathbb{Z}_{n}^{*}$ has an element $a$ of order $m$. An element of order $\lambda(n)$ is often called a primitive $\lambda$-root modulo $n$. We later use two well-known properties of the $\lambda$-function and primitive $\lambda$-roots modulo $n$. The first is a recursive formula for $\lambda(n)$.
[…]
The second property follows from the recursive formula and standard properties of the rings $\mathbb{Z}_{n}^{*}$ (see for example [32]).
Lemma 22.
Let $n,d\in\mathbb{N}$ with $d$ a divisor of $n$. If $a$ is a primitive $\lambda$-root modulo $n$, then $a\mod d$ is a primitive $\lambda$-root modulo $d$.
Finally, by the prime number theorem for arithmetic progressions (see for example [33]) , for every $n$ there also exists a prime $q$ that is a primitive $\lambda$-root modulo $n$. These theorems and facts together lead to the following theorem that will be used for the construction of symplectic lattices and GKP codes from M-SIS lattices.
Theorem 23.
Let $n,q\in\mathbb{N}$ with $q$ a prime that is a primitive $\lambda$-root modulo $2n$. Then the polynomial $X^{n}+1$ modulo $q$ factors as
$X^{n}+1=\prod_{\begin{subarray}{c}d\mid 2n\\ d\nmid n\end{subarray}}\prod_{j=1}^{\phi(d)/\lambda(d)}Q_{d,j}(X),$
where the $Q_{d,j}(X)$ are the irreducible factors of $Q_{d}(X)$ modulo $q$, and each $Q_{d,j}(X)$ has degree $\lambda(d)$.
If prime $q$ is not a primitive $\lambda$-root modulo $2n$, the theorem above can easily be adapted to reflect this, i. e. for each $d$ the number of terms and their degrees have to be adapted according to Theorem 21.
8.2 Symplectic lattices from R-SIS lattices
From now on let $n,q\in\mathbb{N}$ and set $\Phi(X)=X^{n}+1$. We define rings $R,R_{q}$ as above. Let $h(X)=\sum_{i=0}^{n-1}h_{i}X^{i}$ be an element in $R:=\mathbb{Z}[X]/(\Phi(X))$. Consider the linear map $\rho(\textbf{h})$ and its matrix representation with respect to basis $1,X,\ldots,X^{n-1}$ as defined above (see (9) and 10)). Matrix $\rho(\textbf{h})$ is not symmetric. Consequently, matrix $M_{\rho(\textbf{h})}$ as in (3) is not $q$-symplectic. However, define the following (n1)(n-1)-dimensional and $n$-dimensional matrices
$\bar{I}_{n-1}:=\left(\begin{matrix}0&0&\ldots&0&1\\ 0&0&\ldots&1&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&1&\ldots&0&0\\ 1&0&\vdots&0&0\end{matrix}\right),\quad\sigma_{n}:=\left(\begin{matrix}1&0\\ 0_{n}&-\bar{I}_{n-1}\end{matrix}\right).$
Then the matrix
$\bar{\rho}(\textbf{h}):=\sigma_{n}\cdot\rho(\textbf{h}):=\left(\begin{matrix}h_{0}&h_{1}&h_{2}&\ldots&h_{n-1}\\ h_{1}&h_{2}&h_{3}&\ldots&-h_{0}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ h_{n-2}&h_{n-1}&-h_{0}&\vdots&-h_{n-3}\\ h_{n-1}&-h_{0}&-h_{1}&\ldots&-h_{n-2}\end{matrix}\right)$
is symmetric. Hence, it leads to a $q$-symplectic matrix $M_{\bar{\rho}(\textbf{h})}$ as in (3). To simplify notation, we set
$M_{\textbf{h}}:=M_{\rho(\textbf{h})}\quad\text{and}\quad\bar{M}_{\textbf{h}}:=M_{\bar{\rho}(\textbf{h})}.$
In Section 4 we analyze the first successive minimum of ${\cal L}(M_{\bar{\rho}(\textbf{h})})$ for random h. To so we use the following result from Conrad et al. [18] (Lemma 3 and Corollary 1).
Lemma 24.
Let $\textbf{h},M_{\textbf{h}},\bar{M}_{\textbf{h}}$ be as above. Then the lattices ${\cal L}(M_{\textbf{h}})$ and ${\cal L}(\bar{M}_{\textbf{h}})$ are isomorphic. In particular, their first successive minima are identical.
Proof.
Consider the $2n\times 2n$ matrix $U=\sigma_{n}\oplus I_{n}$. Then $U^{T}=U$ and $U^{-1}=U$. Hence,
  1. 1.
    $U$ is orthogonal,
  2. 2.
    $U$ is unimodular.
Furthermore, $U\cdot M_{\textbf{h}}\cdot U=\bar{M}_{\textbf{h}}$. This shows the first statement. The second statement follows from the fact that isomorphic lattices have the same successive minima. ∎
Corollary 25.
Let $\textbf{h},\rho(\textbf{h}),M_{\textbf{h}}$ be as above. Then there exists a $q$-symplectic lattice isomorphic to ${\cal L}(M_{\textbf{h}})$.
Next, we describe ${\cal L}(M_{\rho(\textbf{h})})$ as a module lattice. Later, this allows us to analyze its minimal distance, and, by consequence of Lemma 24, the minimal distance of $q$-symplectic lattice ${\cal L}(M_{\bar{\rho}(\textbf{h})})$. Fix $\textbf{h}\in R$ and $q\in\mathbb{N}$. Set
$\textbf{A}(\textbf{h}):=\left(\begin{matrix}\textbf{h}&-\textbf{1}\end{matrix}\right)\in R^{1\times 2}$
Consider ${\cal L}^{\perp}(\textbf{A}(\textbf{h}))$. Assume that $\textbf{h}\in R_{q}^{\times}$, i.e. h. Viewed as an $R$-module, this lattice has as generating set, in fact as basis, the rows of matrix
$G_{\textbf{h}}:=\left(\begin{matrix}\textbf{1}&\textbf{h}\\ \textbf{0}&\textbf{q}\end{matrix}\right).$
Consequently, viewed as a $\mathbb{Z}$-module, or a lattice, ${\cal L}^{\perp}(\textbf{A}(\textbf{h}))$ has a generating set the rows of matrix $M_{\textbf{h}}$, that is
${\cal L}^{\perp}(\textbf{A}(\textbf{h}))={\cal L}(M_{\textbf{h}})$ (14)
Simplifying notation, we set
${\cal L}^{\perp}(\textbf{h}):={\cal L}(M_{\textbf{h}}).$
Since ${\cal L}(M_{\textbf{h}})$ is isomorphic to the ${\cal L}(\bar{M}_{\textbf{h}})$, lattice ${\cal L}^{\perp}_{\textbf{h}}$ is isomorphic to a $q$-symplectic lattice, namely lattice ${\cal L}(\bar{M}_{\textbf{h}})$. Accordingly, $\frac{1}{\sqrt{q}}{\cal L}^{\perp}_{\textbf{h}}$ is isomorphic to the symplectic lattice $\frac{1}{\sqrt{q}}{\cal L}(\bar{M}_{\textbf{h}})$. By abuse of notation, we call both, $\frac{1}{\sqrt{q}}{\cal L}^{\perp}_{\textbf{h}}$ and $\frac{1}{\sqrt{q}}{\cal L}(\bar{M}_{\textbf{h}})$, symplectic R-SIS lattices. To construct symplectic R-SIS lattices with large minimal distance, we show that for random $\textbf{h}\in\mathbb{Z}_{q}[X]$, lattice $\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{h})$ has large minimal distance. For this, it turns out, the $R$-module view of ${\cal L}^{\perp}(\textbf{h})$ is more convenient.
8.3 Symplectic lattices from M-SIS lattices
We generalize the construction from above to matrices over $R$. The situation studied above, corresponds to matrices with a single row. First some notation. Let $\textbf{H}=(\textbf{h}_{ij})_{1\leq i,j\leq k}$ be a $k\times k$ matrix over $R$. Then we denote by $\rho(\textbf{H})$ the following $nk\times nk$ over $\mathbb{Z}$
$\rho(\textbf{H}):=\left(A(\textbf{h}_{ij})\right)_{1\leq i\leq k},$
that is, we replace entries $\textbf{h}_{ij}$ by the matrix representations of multiplication with $\textbf{h}_{ij}.$ Consider a symmetric matrix H over $R$. Matrix $M_{\rho(\textbf{H})}$ is not symmetric, since matrix representations of multiplications with ring elements are not symmetric. However, for symmetric H the matrix
$\bar{\rho}(\textbf{H}):=\left(\sigma_{n}\cdot A(\textbf{h}_{ij})\right)_{1\leq i\leq k}$
is symmetric.
Further, generalizing the arguments from above, consider the $2nk\times 2nk$ matrix
$\textbf{U}:=\underbrace{\sigma_{n}\oplus\cdots\oplus\sigma_{n}}_{\text{$k$ times}}\oplus\underbrace{I_{n}\oplus\cdots\oplus I_{n}}_{\text{$k$ times}},$
which generalizes the definition above from $k=1$ to arbitrary $k.$ Simple calculations show that
  1. 1.
    U is orthogonal
  2. 2.
    U is unimodular
  3. 3.
    $\textbf{U}\cdot M_{\rho(\textbf{H})}\cdot\textbf{U}=M_{\bar{\rho}(\textbf{H})}.$
Generalizing the simplifying notation from before, we set
$M_{\textbf{H}}:=M_{\rho(\textbf{H})}\quad\text{and}\quad\bar{M}_{\textbf{H}}:=M_{\bar{\rho}(\textbf{H})}.$
This leads to the following generalization of Lemma 24 to symmetric matrices.
Lemma 26.
Let $\textbf{H}\in R^{k\times k}$ symmetric. Define $M_{\textbf{H}},\bar{M}_{\textbf{H}}$ be as above. Then the lattices ${\cal L}(M_{\textbf{H}})$ and ${\cal L}(\bar{M}_{\textbf{H}})$ are isomorphic. In particular, their first successive minima are identical.
As before this yields
Corollary 27.
Let $\textbf{H},\rho(\textbf{h}),M_{\textbf{H}}$ be as above. Then there exists a $q$-symplectic lattice isomorphic to ${\cal L}(M_{\textbf{H}})$
Generalizing further, set
$\textbf{A}(\textbf{H}):=\left(\begin{matrix}\textbf{H}&-\textbf{I}_{k}\end{matrix}\right)\in R^{k\times 2k}$
Consider ${\cal L}^{\perp}(\textbf{A}(\textbf{H}))$. Viewed as an $R$-module, this lattice has as generating set, in fact as basis, the rows of matrix
$G_{\textbf{H}}:=\left(\begin{matrix}\textbf{I}_{k}&\textbf{H}\\ \textbf{0}_{k}&q\textbf{I}_{k}\end{matrix}\right).$
Consequently, viewed as a $\mathbb{Z}$-module, or a lattice, ${\cal L}^{\perp}(\textbf{A}(\textbf{H}))$ has as basis the rows of matrix $M_{\textbf{H}}$, that is
${\cal L}^{\perp}(\textbf{A}(\textbf{H}))={\cal L}(M_{\textbf{H}}).$ (15)
Simplifying notation, we define
${\cal L}^{\perp}(\textbf{H}):={\cal L}^{\perp}(\textbf{A}(\textbf{H})).$
By Lemma 26, $\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H})$ is isomorphic to the lattice $\frac{1}{\sqrt{q}}{\cal L}(\bar{M}_{\textbf{H}})$, which by construction is symplectic. As in the R-SIS case we abuse notation and call $\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H})$ a symplectic M-SIS lattice. To construct symplectic M-SIS lattices with large minimal distance, we show that for random symmetric $\textbf{H}\in\mathbb{Z}_{q}[X]^{k\times m}$, lattice $\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H})$ has large minimal distance.
9 Analysis of minimal distance for symplectic R-SIS and M-SIS lattices
In this section we prove that with probability close to $1$ lattices ${\cal L}^{\perp}(\textbf{h})$ and ${\cal L}^{\perp}(\textbf{H})$ have minimal distance $\approx\sqrt{qn/\pi e}$. Consequently, lattices $\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{h})$ and $\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H})$ have minimal distance $\approx\sqrt{n/\pi e}$ with probability close to $1$. Here the probabilities are with respect to the uniform distribution over elements h in $R_{q}$ and the uniform distribution over symmetric matrices H in $R_{q}^{k\times k},k\in\mathbb{N}.$ The first distribution being the special for $k=1$ of the second distribution. In the sequel we focus on the case of general $k$. But to understand the analysis, it may be useful to think of the special case $k=1$, as this simplifies the notation.
Definition 28.
For $R_{q}$ as above and $k\in\mathbb{N}$ denote the set of symmetric matrices in $R_{q}^{k\times k}$ by $R_{q,k}^{\operatorname{Sym}}$. Hence, $R_{q}=R_{q,1}^{\operatorname{Sym}}$.
The analysis of the minimal distance proceeds in two steps and is mostly identical to the strategy in Section 4. As in this section, it uses standard arguments in lattice-based cryptography (see for example [26]). However, the ring structure of $R$ and $R_{q}$ complicates the analysis significantly. In the first step of the analysis, we determine for fixed $\textbf{z}\in\mathbb{Z}^{2kn},$ interpreted as an element in $R^{2k}$, the probability
$\Pr[\textbf{z}\in{\cal L}^{\perp}(\textbf{H}):\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]$ (16)
This probability depends on the structure of the representative of s in $R_{q}\times R_{q}$ (see Equation 11). For each relevant structure, we derive an upper bound on the number of short z with this structure. Then, the main result on the minimal distance of lattices ${\cal L}^{\perp}(\textbf{H})$ follows by the union bound.
9.1 Analyzing the probability
Let $\Phi(X)=X^{n}+1,R,R_{q}$ as before. Fix an integer $k$. For an element $\textbf{z}\in R^{2k}$, write $\textbf{z}=(\textbf{z}_{1},\textbf{z}_{2}),\textbf{z}_{i}\in R$ and $\textbf{z}_{i}=(\textbf{z}_{i1},\ldots,\textbf{z}_{ik}),\textbf{z}_{ij}\in R$. As before, via its coefficient representation, we can identify an element z in $\mathbb{R}^{2k}$ with an element $\textbf{z}\in\mathbb{Z}^{2kn}$. We use the same notation for both viewpoints. For $\textbf{H}\in R_{q}^{k\times k},$
${\cal L}^{\perp}(\textbf{H})=\left\{\textbf{z}\in R^{2k}:\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod qR\right\}.$
To analyze the probability in (16), we need to further distinguish elements in $\mathbb{R}^{2k}.$ Let
$X^{n}+1:=\prod_{j=1}^{\ell}p_{j}(X)$ (17)
be the factorization of $X^{n}+1$ into irreducible factors modulo $q$ (see (11). Later, using (12) and Theorem 21, we consider this factorization in more detail. For now, set ${\cal I}:=\{1,\ldots,\ell\}$ and for $\textbf{z}\in R^{2k}$
$\bar{{\cal I}}_{\textbf{z}}:=\{j\in{\cal I}:\forall h\textbf{z}_{1h}=\textbf{z}_{2h}=0\mod p_{j}(X)\}\quad\text{and}\quad{\cal I}_{\textbf{z}}:={\cal I}\setminus\bar{{\cal I}}_{\textbf{z}}.$ (18)
Set
$d_{\textbf{z}}:=\sum_{j\in{\cal I}_{\textbf{z}}}\deg(p_{j}(X)),\bar{d}_{\textbf{z}}:=n-d_{\textbf{z}}.$
Lemma 29.
Let $\textbf{z}\in R^{2k},\textbf{z}\neq\textbf{0}\mod qR$. Then
$\Pr[\textbf{z}\in{\cal L}^{\perp}(\textbf{H});\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}})=\frac{1}{q^{d_{\textbf{z}}k}}.$
Proof.
$\textbf{z}\in{\cal L}^{\perp}(\textbf{H})\Leftrightarrow\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod qR\Leftrightarrow\forall j\in{\cal I}\;\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod p_{j}(X).$
[…]
$\Pr[\textbf{H}\textbf{z}_{1}=\textbf{z}_{2}\mod p_{j}(X);\textbf{H}\leftarrow\mathbb{F}_{q^{d_{j}}}^{k\times k}\;\text{symmetric}]=\frac{1}{q^{d_{j}}}.$
For distinct elements in ${\cal I}_{\textbf{z}}$ these events are independent. Observing
$d_{\textbf{z}}=\sum_{j\in{\cal I}_{\textbf{z}}}d_{j}$
proves the lemma. ∎
9.2 Bounds on the number of short vectors
Let $\cal T$$\subset{\cal I}$. In this subsection we derive bounds on the number of short elements in $\textbf{z}=(\textbf{z}_{1},\textbf{z}_{2})\in R^{2k}$ satisfying
$\forall j\not\in{\cal T}:\textbf{z}_{1}=\textbf{z}_{2}=0\mod p_{j}(X).$
To state this more precisely, for ${\cal T}\subseteq{\cal I},{\cal T}\neq\emptyset$, set
  • $p_{\cal T}(X):=\prod_{l\in T}p_{j}(X),\quad\bar{p}_{\cal T}(X):=\prod_{l\not\in T}p_{j}(X)$
  • $d_{\cal T}:=\deg(p_{\cal T}(X)),\quad\bar{d}_{\cal T}:=\deg(\bar{p}_{\cal T}(X))$
  • $g_{\cal T}:=\gcd(n,d_{\cal T})$.
  • $Z_{\cal T}:=\left\{\textbf{z}\in R^{2k}:\textbf{z}_{1}=\textbf{z}_{2}=0\mod\bar{p}_{\cal Z}(X)\right\}.$
Lemma 30.
Let $0<r$ be such that $r\sqrt{q}<q/2$. Then
$\left|B_{2kn}(r\sqrt{q})\cap Z_{{\cal T}}\right|\leq\frac{n}{g_{\cal T}}\sigma_{2kd_{\cal T}}\left(\frac{d_{\cal T}}{n}\right)^{kd_{\cal T}}\left(r+\sqrt{\frac{kn}{2q}}\right)^{2kd_{\cal T}}q^{kd_{\cal T}}$ (19)
Proof.
To simplify notation, we fix set $\cal T$} and drop all subscripts with $\cal T$}, i.e. we use $Z,d,\bar{d},$ and $g$ instead of $Z_{\cal T},d_{\cal T},\bar{d}_{\cal T},g_{\cal T}$.
Let $\textbf{z}\in B_{2kn}(r\sqrt{q})\cap Z_{{\cal T}},\textbf{z}=(\textbf{z}_{11},\ldots,\textbf{z}_{1k},\textbf{z}_{21},\ldots,\textbf{z}_{2k})\in R^{2k}$ with $\textbf{z}_{ij}\in R,1\leq i\leq k,j=1,2$. Set
$\textbf{z}_{ij}^{(h)}:=X^{hg}\cdot\textbf{z}_{ij}\mod X^{n}+1,h=0,\ldots,n/g-1$
The coefficient vector of $\textbf{z}_{ij}^{(h)},$ also denoted by $\textbf{z}_{ij}^{(h)},$ is obtained by cyclically shifting the coefficient vector by $hg$ positions and multiplying the first $gh$ coefficients by 1-1. Each $\textbf{z}_{ij}^{(h)}$ can be written as
$\textbf{z}_{ij}^{(h)}=\overline{\textbf{z}}_{ij}^{(h)}X^{\bar{d}}+\underline{\textbf{z}}_{ij}^{(h)},\deg(\overline{\textbf{z}}_{ij}^{(h)})<d=n-\bar{d},\deg(\underline{\textbf{z}}_{ij}^{(h)})<\bar{d},$
for elements $\overline{\textbf{z}}_{ij}^{(h)},\underline{\textbf{z}}_{ij}^{(h)}\in R$. The elements $\textbf{z}_{ij}^{(h)}$ have the following properties.
  1. 1.
    $\|\textbf{z}_{ij}^{(h)}\|^{2}=\|\textbf{z}_{ij}\|^{2}$ for all indices $i,j,h$,
  2. 2.
    up to sign, each coefficient of $\textbf{z}_{ij}$ appears in exactly $d/g$ elements $\overline{\textbf{z}}_{ij}^{(h)}$,
  3. 3.
    modulo $qR$, each element $\textbf{z}_{ij}^{(h)}$ is uniquely determined by $\overline{\textbf{z}}_{ij}^{(h)}$.
Properties 1. and 2. are immediate. To verify property 3., note that $X^{\bar{d}}$ is invertible modulo all $p_{l}(X)$. Hence, $\textbf{z}_{ij}=0\mod p_{l}(X)$ implies $\textbf{z}_{ij}^{(h)}=0\mod p_{l}(X)$ for all $l\not\in T$. Moreover,
$\underline{\textbf{z}}_{ij}^{(h)}=X^{-\bar{d}}\overline{\textbf{z}}_{ij}^{(h)}\mod p_{l}(X),l\not\in{\cal T},$
where $X^{-\bar{d}}$ denotes the multiplicative inverse of $X^{\bar{d}}$ modulo $p_{l}(X)$. Since the $p_{l}(X)$ are pairwise co-prime, their product is $\bar{p}(X)$, and $\deg(\bar{p}(X))=\bar{d}>\deg(\underline{\textbf{z}}_{ij}^{(h)})$, property 3. follows from the Chinese Remainder Theorem.
By assumption $r\sqrt{q}<q/2$ and, therefore, every element $\textbf{b}\in B_{2kn}(r\sqrt{q})$ must consist of polynomials in $R$, whose coefficients lie between q/2-q/2 and $q/2$. That is, we only need to consider elements in $Z\cap R_{q}^{2k}$ with overall length bounded by $r\sqrt{q}$. Here we take the integers between q/2-q/2 and $q/2$ as a set of representatives for $\mathbb{Z}_{q}$. Next by property 1. and 2.
$\sum_{h=0}^{n/g-1}\sum_{i=1}^{k}\sum_{j=1}^{2}\|\overline{\textbf{z}}_{ij}^{(h)}\|^{2}<\frac{d}{g}r^{2}q.$
Hence, there exists a least one index $t\in\{0,\ldots,n/g-1\}$ such that
$\sum_{i=1}^{k}\sum_{j=1}^{2}\|\bar{\textbf{z}}_{ij}^{(t)}\|^{2}<\frac{d}{n}r^{2}q.$
By property 3. elements $\bar{\textbf{z}}_{ij}^{(t)},1\leq i\leq k,j=1,2$ uniquely determine z and setting
$\bar{\textbf{z}}^{(h)}=(\bar{\textbf{z}}_{11}^{(h)},\ldots,\bar{\textbf{z}}_{1k}^{(h)},\bar{\textbf{z}}_{21}^{(h)},\ldots,\bar{\textbf{z}}_{2k}^{(h)}),$
for every $\textbf{z}\in Z\cap B_{2kn}(r\sqrt{q})$ there is an $t,0\leq t\leq n/g-1,$ and an element $\bar{\textbf{z}}^{(t)}\in B_{2dk}(r\sqrt{qd/n})$ that uniquely determines it.
Therefore
$|Z\cap B_{2kn}(r\sqrt{q})|\leq\frac{n}{g}|\mathbb{Z}^{2dk}\cap B_{2dk}(r\sqrt{qd/n})|.$
By Lemma 7,
$|\mathbb{Z}^{2dk}\cap B_{2dk}(r\sqrt{qd/n})|$ $\leq\sigma_{2dk}(r\sqrt{qd/n}+\sqrt{2dk}/2)^{2dk}$
$=\sigma_{2kd}\left(\frac{d}{n}\right)^{kd}\left(r+\sqrt{\frac{kn}{2q}}\right)^{2kd}q^{kd}$
This proves the lemma. ∎
In some special cases, we can do better than Lemma 30, i.e. we can avoid the term $n/g_{\cal T}$. The next lemma considers the case where $n$ is even and $d=2^{e},e\geq 0,$ is a divisor of $n$, which implies that the polynomial $X^{d}+1$ divides $X^{n}+1$
Lemma 31.
Let $r>0$ be such that $r\sqrt{q}<q/2$. Let $n=2^{e}n^{\prime},e\geq 1,n^{\prime}$ odd. Set $d=2^{e}$ and $p_{\cal T}(X)=X^{d}+1$ for some appropriate subset $\cal T$} of the indices of the divisors of $X^{n}+1$ modulo $q$ (see Equation (17)) . Then
$\left|B_{2kn}(r\sqrt{q})\cap Z_{{\cal T}}\right|\leq\sigma_{2kd}\left(\frac{d}{n}\right)^{dk}\left(r+\sqrt{\frac{kn}{2q}}\right)^{2kd}q^{kd}.$ (20)
Proof.
By assumption
$X^{n}+1=(X^{d}+1)(X^{d(n^{\prime}-1)}-X^{d(n^{\prime}-2)}+\cdots-X^{d}+1),$
i.e. in this case
$\bar{p}_{\cal T}(X)=X^{d(n^{\prime}-1)}-X^{d(n^{\prime}-2)}+\cdots-X^{d}+1.$
Let $\textbf{z}=(\textbf{z}_{11},\ldots,\textbf{z}_{1k},\textbf{z}_{21},\ldots,\textbf{z}_{2k})\in B_{2kn}(r\sqrt{q})\cap Z_{{\cal T}}$. Then there are elements $\textbf{a}_{ij}\in R,1\leq i\leq 2,1\leq j\leq k$, such that as a polynomial element $\textbf{a}_{ij}$ has degree at most d1d-1 and for all index pairs $(i,j)$ we have
$\textbf{z}_{ij}=\textbf{a}_{ij}\bar{p}_{\cal T}\mod qR.$
Because of the special structure of $\bar{p}_{\cal T}(X)$ we get
$\textbf{a}_{ij}\bar{p}_{\cal T}(X)=\textbf{a}_{ij}X^{d(n^{\prime}-1)}-\textbf{a}_{ij}X^{d(n^{\prime}-2)}+\cdots-\textbf{a}_{ij}X^{d}+\textbf{a}_{ij}\mod qR.$
Since $r\sqrt{q}<q/2$, and as in the proof of Lemma 30, we can ignore the modulo $qR$ and restrict ourselves to elements $\textbf{a}_{ij}$ with coefficients less than $r\sqrt{q}<q/2$. Then from the above we conclude that for $\textbf{z}\in B_{2kn}(r\sqrt{q})\cap Z_{{\cal T}}$ we need that
$\|\textbf{a}\|=\left(\sum_{i,j}\|\textbf{a}_{ij}\|^{2}\right)^{1/2}\leq\sqrt{\frac{d}{n}}r\sqrt{q}.$
Since $\textbf{a}\in\mathbb{Z}^{2kd}$ (as a vector), the lemma follows from Lemma 7. ∎
Theorem 32.
With the notation from above
$\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\leq\\ \sum_{{\cal T}\subseteq{\cal I},{\cal T}\neq\emptyset}\frac{n}{g_{\cal T}}\sigma_{2kd_{\cal T}}\left(\frac{d_{\cal T}}{n}\right)^{kd_{\cal T}}\left(r+\sqrt{\frac{kn}{2q}}\right)^{2kd_{\cal T}}.$ (21)
To simplify the notation, denote the right-hand side of the above inequality by
$\epsilon(n,q,k):=\sum_{{\cal T}\subseteq{\cal I},{\cal T}\neq\emptyset}\frac{n}{g_{\cal T}}\sigma_{2kd_{\cal T}}\left(\frac{d_{\cal T}}{n}\right)^{kd_{\cal T}}\left(r+\sqrt{\frac{kn}{2q}}\right)^{2kd_{\cal T}}.$
Proof.
The proof is a simple combination of Lemma 29 and Lemma 30. To simplify notation, define event
$E:=\{\textbf{H}\in R_{q,k}^{\operatorname{Sym}}:\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q}\},$
and denote by $\bar{E}$ its complimentary event. Hence, it suffices to prove that
$\Pr[E(\textbf{H};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\$
By the union bound
$\begin{split}\Pr[\textbf{H}\in E;\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]&=\Pr[\exists\textbf{s}\in B_{2n}(r\sqrt{q})\cap{\cal L}^{\perp}(\textbf{H})\setminus\{0\};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\\ &\leq\sum_{\textbf{s}\in B_{2kn}(r\sqrt{q})\cap\mathbb{Z}^{2kn}\setminus\{0\}}\Pr[\textbf{s}\in{\cal L}^{\perp}(\textbf{H});\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}].\end{split}$
We split the sum into subsets corresponding to subsets $\cal T$$\subset{\cal I}$.
$\sum_{\textbf{s}\in B_{2kn}(r\sqrt{q})\cap\mathbb{Z}^{2kn}\setminus\{0\}}\Pr[\textbf{s}\in{\cal L}^{\perp}(\textbf{H});\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]=\\ \sum_{Z_{\cal T}\subset{\cal I}}\sum_{\textbf{s}\in Z_{\cal T}\cap B_{2kn}(r\sqrt{q})}\Pr[\textbf{s}\in{\cal L}^{\perp}(\textbf{H});\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}].$ (22)
By Lemma 29, for $\textbf{s}\in Z_{\cal T}$,
$\Pr[\textbf{s}\in{\cal L}^{\perp}(\textbf{H});\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\leq\frac{1}{q^{kd_{\cal T}}},$
and by Lemma 30, for $\cal T$$\subset{\cal I}$,
$|Z_{\cal T}\cap B_{2kn}(r\sqrt{q})|\leq\frac{n}{g_{\cal T}}\sigma_{2kd_{\cal T}}\left(\frac{d_{\cal T}}{n}\right)^{kd_{\cal T}}\left(r+\sqrt{\frac{kn}{2q}}\right)^{2kd_{\cal T}}q^{kd_{\cal T}}.$
Together with (22) this proves the theorem. ∎
Corollary 33.
With the notation from above
$\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\leq\\ \sum_{{\cal T}\subseteq{\cal I},{\cal T}\neq\emptyset}\frac{n}{g_{\cal T}}\frac{1}{\sqrt{2\pi d_{\cal T}k}}\left(\frac{\pi e}{nk}\right)^{kd_{\cal T}}\left(r+\sqrt{\frac{kn}{2q}}\right)^{2kd_{\cal T}}.$ (23)
Proof.
For every integer $d$, $\sigma_{2d}=\frac{\pi^{d}}{d!}$. Using Stirling’s formula (Lemma 8), we obtain for every set $\cal T$}
$\sigma_{2kd_{\cal T}}\left(\frac{d_{\cal T}}{n}\right)^{kd_{\cal T}}\leq\frac{1}{\sqrt{2\pi d_{\cal T}k}}\left(\frac{\pi e}{d_{\cal T}k}\right)^{kd_{\cal T}}.$
This proves the corollary ∎
Remark 34.
Using Lemma 31, in Theorem 32 as well as in Corollary 33, for all subsets $\cal T$} such that $p_{\cal T}(X)$ is of the form $X^{2^{e}}+1,e\geq 0$, we can drop the factor $n/g_{\cal T}$.
10 Main results - module and ring constructions
In this section we use Theorem 32 and Corollary 33 in combination with Theorem 23 to construct $n$-dimensional symplectic lattices with minimal distance $\approx\sqrt{n/\pi e}$, and then, as in Section 5, good GKP codes. Theorem  32 and Corollary 33 only yield meaningful results, if the right-hand sides in equations ( 21) and ( 23) do not contain to many terms. Equivalently, the factorization in Theorem 23 may not contain too many terms. When $X^{n}+1$ factors into a product of linear terms, this is certainly not the case [^4] Note that this case is interesting in many cryptographic applications of R-SIS lattices. Since then the number of terms in (21) in exponential in $n$. Fortunately, Theorem 23, together with (13), offers us interesting and useful alternatives. More precisely, the results we present below concentrate on choices of $n,q,r$ such that
  1. 1.
    the number of terms in the sums of the right-hand sides of (21) and  (23) are small,
  2. 2.
    for all subsets $\cal T$}, except possibly one, the degrees $d_{\cal T}$ are large, i. e. $d_{\cal T}=\Omega(n^{\delta}),\delta>$, and for the exceptional set $\cal T$} the degree $d_{\cal T}$ is a power of $2$.
We consider the case $n=2^{e}p_{1}^{e_{1}}p_{2}^{e_{2}},$ where $p_{1},p_{2}$ are distinct odd primes. Then, according to Theorem 12, the factor of $X^{n}+1$ in $\mathbb{Z}$ are $\Phi_{2d}(X),$ where
$2d\in\{2^{e+1}p_{1}^{d_{1}}p_{2}^{d_{2}}\mid 0\leq d_{1}\leq e_{1},0\leq d_{2}\leq e_{2}\}.$
To apply Theorem 23 with some primitive $\lambda$-root $q$ of $2n$, and hence primitive $\lambda$-root for all $2d,d|n$, we need determine the values $\phi(2d),\lambda(2d),$ and $\phi(2d)/\lambda(2d),$. To do so, for $2d=2^{e+1}p_{1}^{d_{1}}p_{2}^{d_{2}}$, write
$\phi(p_{i}^{d_{i}})=p_{i}^{d_{i}-1}(p_{i}-1))=2^{f_{i}}u_{i}p_{i}^{d_{i}-1},\;\text{$f_{i}\geq 1,u_{i}$ odd},$
and set $u=\gcd(p_{1}^{d_{1}-1}u_{1},p_{2}^{d_{2}-1}u_{2})$. Then, by straightforward calculations,
$\lambda(d)=\begin{cases}2^{\max\{e,f_{1},f_{2}\}}\frac{\phi(p_{1}^{d_{1}}\phi(p_{2}^{d_{2}})}{2^{f_{1}+f_{2}}u}&e=0,1\\ 2^{\max\{e-1,f_{1},f_{2}\}}\frac{\phi(p_{1}^{d_{1}}\phi(p_{2}^{d_{2}})}{2^{f_{1}+f_{2}}u}&e\geq 2.\end{cases}$ (24)
Hence
$\frac{\phi(2d)}{\lambda(2d)}=\begin{cases}2^{e+f_{1}+f_{2}-\max\{e,f_{1},f_{2}\}}\cdot u&e=0,1\\ 2^{e+f_{1}+f_{2}-\max\{e-1,f_{1},f_{2}\}}\cdot u&e\geq 2.\end{cases}$ (25)
These formulas also apply if $d_{1}=0$ or $d_{2}=0$. For the two conditions form above to hold we want
  1. 1.
    $e_{1},e_{2}$ to be small (this limits the number of factors $\Phi_{2d}(X)$ of $X^{n}+1$ over $\mathbb{Z}$)
  2. 2.
    for all $2d$, the values $f_{1},f_{2},u$ defined above to be small (this limits the number of factors of $\Phi_{2d}(X)$ modulo $q$).
The first, and most important, of our constructions construction uses $n=2^{e}$. This setting is also by Conrad et al. in  [18]. But whereas Conrad et al. use NTRU lattices we use M-SIS lattices from cryptography. Other possible choices for $n,q$ are discussed below. To choose appropriate values for $r$ and analyze the upper bounds for the probability of small minimal distance in Theorem 32 and Corollary 33, we concentrate on subsets $\cal T$} leading to polynomials with smallest degree $d:=d_{\cal T}$ and consider the term
$\operatorname{smd}(n,k,q,r,d):=\frac{1}{\sqrt{2\pi kd}}\left(\frac{\pi e}{nk}\right)^{kd}\left(r+\sqrt{\frac{nk}{2q}}\right)^{2kd}.$ (26)
Setting
$\rho=\rho(n,k,\gamma):=r+\sqrt{\frac{nk}{2q}}=e^{-\gamma/2}\sqrt{\frac{nk}{\pi e}},\;\text{for some}\;\gamma\in\mathbb{R},$ (27)
we get
$\operatorname{smd}(n,k,q,r,d)=\frac{1}{\sqrt{2\pi dk}}e^{-\gamma dk}\;\text{and}\;r\geq\sqrt{\frac{nk}{\pi e}}\left(e^{-\gamma/2}-\sqrt{\frac{\pi e}{2q}}\right).$ (28)
Finally, set
$q\geq\frac{2\pi e}{\gamma^{2}}.$ (29)
Using, $e^{x}>1+x$ for all $x$, we get
$r>\begin{cases}\sqrt{\frac{nk}{\pi e}}(1-\gamma),&\gamma\geq 0\\ \sqrt{\frac{nk}{\pi e}},&\gamma<0.\end{cases}$ (30)
With these settings and bounds first we get the following results.
Theorem 35.
Let $k,n,q\in\mathbb{N}$ with $n$ a power of $2$ and $q$ a prime with $q=5\mod 8$ or $q=3\mod 16$. Set $R:=\mathbb{Z}[X]/(X^{n}+1)$ and $R_{q}:=\mathbb{Z}[q]/(X^{n}+1)$. Choose H uniformly at random from $R_{q,k}^{\operatorname{Sym}}$ and set
${\cal L}:=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H}).$
  1. 1.
    For $q\geq\frac{\pi e(nk)^{2}}{2}$, with probability at least $1-\frac{e^{2}+2\sqrt{2}e}{\sqrt{2\pi nk}}=1-\Omega(1/\sqrt{nk})$
    $\lambda_{1}({\cal L})\geq\sqrt{\frac{nk}{\pi e}},$
  2. 2.
    For $q\geq 2\pi e(nk)^{3/2}$, with probability at least $1-e^{-(nk)^{1/4}}$
    $\lambda_{1}({\cal L})\geq\left(1-\frac{1}{(nk)^{3/4}}\right)\sqrt{\frac{nk}{\pi e}}.$
Proof.
The proof of Theorem 2’ in Chapter 4 of [32] shows that for $q=5\mod 8,n\geq 8$, the order of $q$ modulo $2n$ is $n/2$. That is, $q$ is a primitive $\lambda$-root modulo $n$. The same proof can be used for $q=3\mod n$ and $n\geq 16$.
By Theorem 23, Equation (25) and with our choice of $q$, modulo $q$ polynomial $X^{n}+1$ factors into two polynomials $p_{1}(X),p_{2}(X)$ of degree $n/2$ each. Hence the right-hand side in (23) has three summands for the sets ${\cal T}=\{1\},\{2\},\{1,2\}$. The corresponding degrees $d_{\cal T}$ are given by $n/2,n/2,n$ with values $g_{\cal T}$ and $n/g_{\cal T}$ given by $n/2,n/2,n$ and $2,2,1,$ respectively. Using Corollary 33 and the notation and bounds from above, we obtain
$\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]\leq\\ 2\cdot\operatorname{smd}(n,k,q,r,n/2)+\operatorname{smd}(n,k,q,r,n).$ (31)
  1. 1.
    In this case, using the notation from above, we set $\gamma:=-2/kn$ and, correspondingly, choose a prime $q\geq\pi e(kn)^{2}/2$ (see (29)). This yields
[…]
    and
[…]
  2. 2.
    Here we choose
    $\gamma:=\frac{2}{(nk)^{3/4}}\;\text{and}\;q\geq\frac{\pi e(nk)^{3/2}}{2},$
    and use the same line as arguments as in the previous case.
Next we consider the case where $n$ is the product of $2$ and two odd primes.
Theorem 36.
[…]
${\cal L}:=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H}).$
  1. 1.
    For $k>\ln(n/2)/4$ and $q\geq\frac{32\pi ek^{2}}{\ln(n/2)^{2}}$, with probability at least $1-\frac{1+e^{-\Omega(p_{1})}}{\sqrt{2\pi nk}}=1-\Omega(1/\sqrt{nk})$
    $\lambda_{1}({\cal L})\geq\left(1-\frac{\ln(n/2)}{4k}\right)\sqrt{\frac{nk}{\pi e}}.$
  2. 2.
    For $4k>n^{1/4}$ and $q\geq\frac{32\pi ek^{2}}{\sqrt{n}}$, with probability at least $1-(e^{-p_{1}^{1/4}}+e^{-\Omega(p_{1})}$
    $\lambda_{1}({\cal L})\geq\left(1-\frac{n^{1/4}}{4k}\right)\sqrt{\frac{nk}{\pi e}}.$
Proof.
From Theorem 23 and Equations (24), (25) we get that modulo $q$ the polynomial $X^{n}+1$ factors into the polynomial $X^{2}+1$, two polynomials of degree $p_{1}-1$ (factors of $\Phi_{4p_{1}}(X)$ modulo $q$), two polynomials of degree $p_{2}-1$ (factors of $\Phi_{4p_{1}}(X)$ modulo $q$), and four polynomials of degree $\frac{(p_{1}-1)(p_{2}-1)}{2}$ (factors of $\Phi_{4p_{1}p_{2}}(X)$ modulo $q$). Hence, the right-hand sum in Equation 23 consists of $511$ terms. The smallest degree $d_{\cal T}$ is $2$ and the next smallest degree is $p_{1}-1$. The (asymptoktically) determining term corresponds to the polynomials $X^{2}+1$. In which case we can use Remark 34 and focus on $\operatorname{smd}(n,k,q,r,2)$.
  1. 1.
    We choose
    $\gamma:=\frac{\ln(n/2)}{4k}\quad\text{and}\quad q\geq\frac{\pi e32k^{2}}{\ln(n/2)}.$
    As in the proof of the previous theorem, we can show that with the choices
    $\operatorname{smd}(n,k,q,r,2)\leq\frac{1}{2\pi nk}\quad\text{and}\quad r\geq\left(1-\frac{\ln(n/2)}{4k}\right)\sqrt{\frac{nk}{\pi e}}.$
    For all other sets $\cal T$} with corresponding degree $d=d_{\cal T}$ we have $d\geq p_{1}-1$ and hence
    $\operatorname{smd}(n,k,q,r,d)<e^{-\gamma kd}\leq e^{-p_{1}/8}.$
    This implies that for the other set $\cal T$} their combined contribution to $\Pr[\lambda_{1}({\cal L}^{\perp}(\textbf{H}))<r\sqrt{q};\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}]$ is bounded by
    $511n\operatorname{smd}(n,k,q,r,d)=e^{-\Omega(p_{1})}.$
    This proves the first part of the theorem.
  2. 2.
    The proof of the second part of the theorem follows as the proof for the first part by choosing
    $\gamma=\frac{n^{1/4}}{4k}\quad\text{and}\quad q\geq\frac{32\pi ek^{2}}{n^{1/2}}.$
Finally, consider the case $n=2^{e}p,e\geq 2,p$ prime. We only state the result. Its proof follows along the same lines as the previous two proofs.
Theorem 37.
Let $k,n,q\in\mathbb{N}$ with $n=2^{e}p,e\geq 2$, and $p,q$ primes such that $q$ is a primitive $\lambda$-root modulo $2n$. Set $R:=\mathbb{Z}[X]/(X^{n}+1)$ and $R_{q}:=\mathbb{Z}[q]/(X^{n}+1)$. Choose H uniformly at random from $R_{q,k}^{\operatorname{Sym}}$ and set
${\cal L}:=\frac{1}{\sqrt{q}}{\cal L}^{\perp}(\textbf{H}).$
  1. 1.
    For $k>\ln(n/2^{e})/2^{e+1}$ and $q\geq\frac{2^{2e+3}\pi ek^{2}}{\ln(n/2^{e})^{2}}$, with probability at least $1-\frac{1+e^{-\Omega(n)}}{\sqrt{2\pi nk}}=1-\Omega(1/\sqrt{nk})$
    $\lambda_{1}({\cal L})\geq\left(1-\frac{\ln(n/2^{e})}{2^{e+1}k}\right)\sqrt{\frac{nk}{\pi e}}.$
  2. 2.
    For $2^{e+1}k>n^{1/4}$ and $q\geq\frac{2^{2e+3}\pi ek^{2}}{\sqrt{n}}$, with probability at least $1-(e^{-n^{1/4}}+e^{-\Omega(n)})$
    $\lambda_{1}({\cal L})\geq\left(1-\frac{n^{1/4}}{2^{e+1}k}\right)\sqrt{\frac{nk}{\pi e}}.$
[…]
11 Bounded distance decoding in the module case and the role of $k$
We can also use algorithm $\textsc{Bdd}_{\text{triv}}$ from Section 6 to decode the GKP codes obtained from M-SIS lattices. The most time consuming step in algorithm $\textsc{Bdd}_{\text{triv}}$ is the matrix-vector multiplication c’H in step 2, which requires $n^{2}$ multiplications in $\mathbb{F}_{q}$. If $n$ is a power of $2$, instead of the SIS lattice construction of GKP codes we can use the M-SIS-based constructed using Theorem 35, with $k=1$. The matrix-vector multiplication in step $2$ of algorithm $\textsc{Bdd}_{\text{triv}}$ becomes a multiplication of two ring elements in $R=Z[X](X^{n}+1)$. By construction, this ring contains $n$-th roots of unity which allows us to perform multiplication in $R$ with ${\textbf{{O}}}(n\log(n))$ arithmetic operations in $\mathbb{F}_{q}$ using the fast Fourier transform (FFT).
However, powers of $2$ are relatively sparse. Using $k>1$, allows us to obtain fast decoding for more general $n$. If we choose the dimension of the code as $n\cdot k$ with $n$ being a power of $2$, the multiplication c’H requires $k^{2}$ multiplications in ring $R=\mathbb{Z}[X]/(X^{n}+1)$ at the cost ${\textbf{{O}}}(k^{2}n\log(n))$ arithmetic operations in $F_{q}$, again using FFT. Hence, the parameter $k$ and the use of M-SIS lattices and not just R-SIS lattice (corresponding to $k=1$) allows us more flexibility in designing code GKP codes with efficient decoding algorithms. Using other choices for $n$ than just powers of $2$, yields yet more flexibility in designing good GKP codes for various dimensions.
12 Simulation of main results for ring constructions
In this section we show some simulation results in Section 10, i.e. the minimal distance for symplectic R-SIS and M-SIS lattices. The decoding performance for these lattices shows no difference from SIS lattices in Section 7, therefore we don’t put the results here.
(a) $n=4$, $m\in[1,5]$, $q=701$
(b) $n=2^{e}$, $m=1$, $q=3$
(c) $n=2^{e}p_{1}p_{2}$
(d) $n=2^{e}p$
Figure 3: Simulations of minimal distance for symplectic R-SIS and M-SIS lattices. (a) $n=4$, $m\in[1,5]$, $q=701$. (b) $n=2^{e}$, $m=1$, $q=3$. (c) $n=2^{e}p_{1}p_{2}$. (d) $n=2^{e}p$.
Figure 3 shows the simulation results of minimal distance for symplectic R-SIS and M-SIS lattices. For each combination of $(n,k,q)$, 1000 lattice samples are generated with $\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}$. Minimal distance of lattice is then calculated via BKZ reduction. In the figure, the red line represents the asymptotic bound $\sqrt{nk/\pi e}$ and the black dots represents the minimal distance for each lattice sample. The numbers at the bottom shows that for each combination $(n,k)$, how many samples have minimal distance $\lambda_{1}>\sqrt{nk/\pi e}$. Figure 3(a) corresponds to the case of Theorem 35, where $n=4$ is a power of 2, and $q=701$ satisfies $q=5$ mod $8$ and $q\geq\frac{\pi e(4\cdot 5)^{2}}{2}$. As shown in the figure, in all combinations of $(n,k)$, the generated M-SIS lattices have minimal distance larger than $\sqrt{nk/\pi e}$ with probability almost to 1, which is even much larger then what the theorem tells; Figure 3(b) also has $n$ equals a power of 2, but a small $q=3$ is used in the simulation. Still the generated R-SIS lattices have minimal distance larger than $\sqrt{nk/\pi e}$ with probability almost to 1. Thus from the simulation we make the conjecture that the generated M-SIS lattices have a high probability to achieve $\lambda_{1}\geq\sqrt{nk/\pi e}$ regardless of what $q$ is used for their generation. Figure 3(c) corresponds to the case of Theorem 36, where $n$ is a product of 2 primes. And for each $(n,k)$, certain $q$ equal to a primitive $\lambda$-root modulo $2n$ as well as $q\geq\frac{32\pi ek^{2}}{\ln(n/2)^{2}}$ is used in the simulation. The lattices still have higher probability to achieve $\lambda_{1}\geq\sqrt{nk/\pi e}$ than theorem suggested, but the probability is lower than the case where $n$ is a power of 2. This can relate to the factorization of polynomial $X^{n}+1$, which is mentioned in the proof of Theorem 36. Figure 3(d) corresponds to the case of Theorem 37, where $n$ is a product of 4 and a prime. Certain $q$ corresponds to the fist part of the theorem is used in the simulation. Again the results corresponds well with the theorem, and even suggest a higher probability of $\lambda_{1}\geq\sqrt{nk/\pi e}$.
13 Conclusion and Outlook
In this paper, we give the first efficient constructions of good families of GKP codes. Our constructions achieve the code distances non-constructively shown to be achievable by Preskill and Harrington [14] (based on results on symplectic lattices by Buser and Sarnak [19]) and conjectured by Conrad et al. to be achievable constructively by NTRU-based GKP codes [18]. Furthermore, the achieved distances are provably optimal or near optimal (up to small constants). Our GKP codes are constructed from well-known families of cryptographic lattices, SIS lattices. We also showed a trivial decoding algorithm for our GKP codes which does not use trapdoors. Experiments confirm the effectiveness of the constructions. For parameter choices and decoding algorithms the experiments even supersede the theoretical results.
Our results raise various interesting research questions. The NTRU constructions in [18] use trapdoors to obtain improved decoding algorithms for the NTRU-based GKP codes. The SIS lattices that we use in our constructions do not have trapdoors. Hence the questions arises whether we can use other cryptographic lattices with trapdoors to construct good GKP codes with efficient decoding algorithms exploiting the trapdoors. Examples of such lattices include learning with error (LWE) lattices. Also lattices with so called $G$-trapdoors may be of interest (for both see  [25]). Our results for M-SIS lattices, though general, lead to non-trivial results only for limited parameter settings. We conjecture that techniques developed for the ring-LWE and module-LWE problems (see [34]) can be used to obtain improved results applicable to more general parameter settings. Finally, and practically most importantly, GKP codes as presented in this paper need to be combined with other quantum error-correcting codes like surface codes or CSS-codes to obtain efficient codes applicable for the design of large scale quantum computers. Hence, what are the best and most efficient combinations of GKP and other quantum error-correcting codes?
For randomized GKP codes constructed from cryptographic lattices, one can also think about exploiting their randomized and cryptographic property. For example, Conrad et al. [18] proposed a quantum communication protocol with NTRU-based GKP codes. The protocol is said to be secure with the trapdoors used in NTRU systems. Similar protocol should also work for other code constructions that have trapdoors such as LWE. For GKP codes that do not have trapdoors in their construction, it may also be possible to use them for constructing secure quantum communication protocols by using the randomness in the codes. The randomness of these codes and their relations with cryptographic lattices may also lead to possible decoding algorithms which are more efficient or better-performed compared with normal GKP codes.
Acknowledgments
This work was funded by the Ministry of Culture and Science of the State of North Rhine-Westphalia and by the Federal Ministry of Research, Technology and Space (BMBFTR) (project PhoQuant). Additionally, this work was supported by the Equal Opportunity Program, Grant Line 2: Support for Female Junior Professors and Postdocs through Academic Staff Positions, 15th funding round of Paderborn University.